### v1

# RLIE: Rule Generation with Logistic Regression, Iterative Refinement, and Evaluation for Large Language Models

Abstract
Nowadays, Large Language Models (LLMs) are able to propose rules in natural language, overcoming constrains of a predefined predicate space inherent in traditional rule learning. However, existing methods using LLMs often overlook the combination effects of rules, and the potential of coupling LLMs with probabilistic rule learning to ensure robust inference is not fully explored. To address this gap, we introduce RLIE, a unified framework that integrates LLMs with probabilistic modeling to learn a set of probabilistic rules. The RLIE framework comprises four stages: (1) Rule generation, where a LLM proposes and filters candidate rules; (2) Logistic regression, which learns the probabilistic weights of the rules for global selection and calibration; (3) Iterative refinement, which continuously optimizes the rule set based on prediction errors; and (4) Evaluation, which compares the performance of the weighted rule set as a direct classifier against various methods of injecting the rules into an LLM. Generated rules are the evaluated with different inference strategies on multiple real-world datasets. While applying rules directly with corresponding weights brings us superior performance, prompting LLMs with rules, weights and classification results from the logistic model will surprising degrade the performance. This result aligns with the observation that LLMs excel at semantic generation and interpretation but are less reliable at fine-grained, controlled probabilistic integration. Our work investigates the potentials and limitations of using LLMs for inductive reasoning tasks, proposing a unified framework which integrates LLMs with classic probabilistic rule combination methods, paving the way for more reliable neuro-symbolic reasoning systems.
1 Introduction
In data-driven applications and scientific discovery, the goal is not merely to predict outcomes, but to construct a set of verifiable, reusable, and composable theories(Zhou et al., 2024; Yang et al., 2024a; Minh et al., 2022). These theories can enable explainable, auditable decisions while driving the discovery of new knowledge and underlying structures(Yang et al., 2023; 2024b). These theories can be expressed in formal, structural statements(Cohen et al., 1995; Cropper & Morel, 2021) or natural language hypotheses(Zhou et al., 2024), and they share a common characteristic: they are declarative, testable, and self-contained discriminative patterns that yield predictions verifiable by external evidence [^1] In this paper, we do not distinguish between the terms “rule” and “hypothesis”, and will use “rule” throughout the text for consistency.. The challenge of how to learn and integrate a set of such collaborative patterns to both uncover regularities and advance knowledge accumulation has long been a central problem in machine learning and artificial intelligence(Bazgir et al., 2025).
Take the binary classification task of spam detection as an example. A set of rules might include (1) $\mathrm{HasToken}(\text{`win prize'},x)\land\mathrm{HasAttachment}(x)$; (2) $\mathrm{DomainInBlacklist}(x)\land\mathrm{HasShortLink}(x)$; and (3) $\mathrm{TitleAllCaps}(x)\land\mathrm{HasToken}(\text{`urgent'},x)$, all implying the head “is spam”. These rules cover different facets of the problem, and must work in concert. A classic approach to use the rules to do reasoning is deterministic aggregation, where a positive classification is made, if any rule is satisfied(Fürnkranz & Kliegr, 2015). A probabilistic approach, in contrast, treats each rule’s classification result as a binary indicator ($r_{j}\in\{0,1\}$) and employs a statistical model, such as logistic regression, to learn a global set of weights for combining these signals into a unified prediction(Ruczinski et al., 2003; Friedman & Popescu, 2008). While the well-defined structure of rules offers significant advantages, traditional rule learning methods are constrained by a predefined predicate space, which limits their ability to handle unstructured data and open-ended semantics, thereby restricting their applicability in broader scenarios(Cerna & Cropper, 2024; Hocquette et al., 2024).
The recent advent of Large Language Models (LLMs) presents a new opportunity for rule learning. Given a few examples and task context, LLMs can directly generate rules in natural language(Ellis, 2023; Singh et al., 2022). Their linguistic richness and contextual alignment facilitate the articulation of fine-grained conditions, relations and exceptions, covering patterns that are difficult to capture with manually defined predicates. However, existing methods often focus on the iterative optimization of a single rule(Qiu et al., 2023) or the independent generation and simple aggregation of multiple rules(Yang et al., 2023; Zhou et al., 2024; Yang et al., 2024b), failing to jointly learn and compress the rules as a cohesive set. Concurrently, probabilistic approaches in the context of LLMs remains largely unexplored, as LLMs’ ability to robustly learn the rule weights from data and apply them to perform reliable probabilistic reasoning during inference is not systematically examined yet. This gap motivates our pursuit of a framework that combines the generative and expressive capabilities of LLMs with the probabilistic composition and calibration of traditional methods to produce a flexible, robust and reusable rule set.
To achieve this synergy, we propose RLIE, a unified framework for LLM-based rule learning that integrates Rule generation, Logistic regression, Iterative refinement and Evaluation. First, in the rule generation stage, an LLM produces a set of independent, individually assessable natural language rules based on a few examples and the task context. Second, for probabilistic combination, a regularized logistic regression model is employed to perform global weighting and selection on the rules, yielding a compact, interpretable rule set with importance weights. Third, in the iterative refinement stage, challenging examples are identified based on the prediction errors. These examples, along with the current rule set, are fed back to the LLM to prompt reflective generation of improved rules until convergence. Finally, a evaluation of two primary inference strategies is conducted: (1) direct inference using the weighted model, and (2) injecting the rule information into an LLM. For the latter, we systematically examine three levels of information: (i) providing only the rule, (ii) providing the rules with their weights, and (iii) providing the rules, weights and the linear model’s prediction as a reference. The layered design allows for a systematic evaluation of quality of rules and analysis of the trade-offs in accuracy and robustness for different ways of using rules. Using our framework, we can get high quality rules in natural languages, with optimal ways of utilizing them to perform rule based reasoning. The main contribution of this paper are as follows:
  1. 1.
    We propose RLIE, a unified paradigm for LLM-based probabilistic rule learning. It combines the natural language rule generation of LLMs with the global weighting and selection of regularized logistic regression to produce a self-consistent, collaborative and auditable rule set.
  2. 2.
    We design a hierarchical set of inference strategies to systematically compare direct inference with the linear combiner against three methods of injecting information into an LLM: rules-only, rules+weights and rules+weights+linear prediction reference. We provide empirical analysis and practical guidance regrading their respective trade-offs in terms of accuracy, robustness and calibration.
  3. 3.
    On multiple real-world datasets, RLIE achieves superior over all performance compared to a range of LLM-based methods. The learnt rule sets are more compact and semantically clearer, prompting knowledge discovery and human-AI consensus.
2 Related Work
2.1 Classical Rule Learning and Its Applications
Rule Learning. Canonical approaches to rule learning include Inductive Logic Programming (ILP)(Cropper et al., 2022), which is centered on search and constraint optimization; association and pattern mining methods that rely on heuristics and pruning; and differentiable neuro-symbolic methods(Qiao et al., 2021; Glanois et al., 2022; Yang et al., 2024a) that couple symbolic rules with continuous representations. A common objective across these paradigms is to produce a set of human-readable and reusable discriminative rules from a structured candidate space. However, they are fundamentally dependent on a predefined predicate space.
Rule Organization and Reasoning. In the traditional paradigm, the interaction among rules is determined by their organization and the semantics of the inference process. One approach is the use of ordered rule lists or decision lists(Yang et al., 2017; Xu et al., 2024), which employ sequential matching to explicitly encode relationships of priority, exception, and default cases. This is well-suited for scenarios with clear sequential or override logic. Another approach involves unordered rule sets(Qiao et al., 2021; Yang & van Leeuwen, 2022), which support the parallel integration of multiple pieces of evidence, facilitating global trade-offs and model compression. Within the unordered setting, deterministic aggregation methods (e.g., logical OR/AND, thresholding, majority voting) are simple and intuitive but offer limited flexibility when faced with local conflicts, coverage gaps, or inter-rule dependencies. Consequently, many works have introduced probabilistic aggregation to model uncertainty and improve calibration. A common and interpretable implementation is the linear log-odds model, which treats the binary satisfaction of each rule $r_{j}(x)\in\{0,1\}$ on a sample $x$ as a feature and learns a weighted linear combination(Ruczinski et al., 2003):
$\mathrm{Pr}(y=1\mid x)=\sigma(\beta_{0}+\sum_{j}\beta_{j}r_{j}(x)).$
Our work builds upon this foundation. Traditional methods typically assume rules are expressed in a fixed, executable symbolic form, allowing for direct evaluation of their satisfaction. In contrast, our approach focuses on rules expressed in natural language, where satisfaction must be determined by an LLM. We adopt a two-level design: at the “local judgment” level, we leverage the expressive power and adaptability of LLMs to interpret individual rules; at the “global aggregation” level, we retain the robust and calibratable probabilistic linear combiner from the classical paradigm. This hybrid design captures the flexible semantics of natural language rules while ensuring the overall inference process remains transparent and robust.
2.2 LLM-based Rule Learning and Its Applications
Rule Learning. Large Language Models (LLMs) can directly propose hypotheses in natural language. A representative approach involves a “single-hypothesis refinement” loop, where the model maintains only the current best hypothesis and iteratively improves its quality and interpretability through feedback from validation sets or external executors(Qiu et al., 2023). In contrast, other methods adopt a “multi-hypothesis set” approach. For example, HypoGeniC(Zhou et al., 2024) generates multiple candidate hypotheses from data samples, maintains a top-$k$ set, and uses training signals to filter and update this set, ultimately forming an interpretable hypothesis bank. These two strategies exemplify the trade-off between deep optimization of a single hypothesis and the parallel retention of multiple diverse hypotheses.
Rule Organization and Reasoning. Rules expressed in natural language are typically reasoned over by the LLM that generated them. Existing work often passes one or more rules as context to an LLM, demonstrating that such information can enhance its reasoning capabilities(Zhang et al., 2024; Zhou et al., 2024). However, as the rules become more numerous or complex, the stability and instruction-following fidelity of LLMs can become unreliable. The black-box nature of LLMs makes it challenging to analyze the organization and interaction of these rules.
Our work is distinct from existing LLM-based rule learning approaches in that we are the first to explicitly combine LLMs with probabilistic methods to learn a set of weighted rules. This integration enables more principled sample selection during the iterative refinement stage and leads to more flexible and robust inference. To better analyze the synergy between this new rule paradigm and LLMs, we systematically compare direct inference using the probabilistic combiner with a hierarchy of strategies for injecting the rules, their weights, and the combiner’s predictions into an LLM. This allows us to provide empirical analysis and practical guidance on the trade-offs in accuracy and robustness.
3 Method
We consider a binary classification task on a labeled dataset $\mathcal{S}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}\in\mathcal{X}$ is a natural language text and $y_{i}\in\{0,1\}$ is its corresponding label. The dataset $\mathcal{S}$ is partitioned into a training set $\mathcal{S}_{\mathrm{tr}}$, a validation set $\mathcal{S}_{\mathrm{val}}$ and a test set $\mathcal{S}_{\mathrm{te}}$. To perform rule based reasoning, our objective is to learn a set of natural language rules $\mathcal{H}^{\star}=\{h_{1},\ldots,h_{m}\}$ and a set of weights $\theta$ that aggregate the rules to model the discriminative relationship between $x$ and $y$. Here, $m\leq H$ where $H$ is a predefined capacity limit, and $h_{i}$ is a rule in natural language. Our method contains four stages, and the whole procedure is illustrated in Figure 1:
Figure 1: The pipeline of the RLIE framework. (1) Rule Generation: An LLM proposes an initial set of rules from a small sample of training data. (2) Logistic Regression: A regularized logistic regression model is trained on the full training set to learn probabilistic weights for the current rules. (3) Iterative Refinement: Prediction errors from the regression model are used to mine hard examples, which are then used to prompt the LLM to refine or generate new rules. This cycle continues until performance on a validation set converges. (4) Evaluation: The learnt rules as well as weights are evaluated with different inference strategies. Detailed explanations of different strategies are in Section 3.4.
  1. 1.
    Rule Generation: We initialize an empty rule set $\mathcal{H}:=\emptyset$. An LLM is prompted with a randomly selected subset of $\mathcal{S}_{\mathrm{tr}}$ to generate candidate rules, which are then filtered based on coverage to form the initial rule set $\mathcal{H}^{(1)}$.
  2. 2.
    Logistic Regression: To obtain the optimal weights to combinatorially use the generated rules, logistic regression is performed to learn the rule weights $\theta=(\bm{\beta},b)$ on $\mathcal{S}_{\mathrm{tr}}$ with hyperparameters tuned using $\mathcal{S}_{\mathrm{val}}$.
  3. 3.
    Iterative Refinement: Hard examples are selected from $\mathcal{S}_{\mathrm{tr}}$ based on prediction errors. These samples, along with the current rule set $\mathcal{H}^{(t)}$, are fed back to the LLM to generate a set of new rules to revise existing rules. The union of newly generated rules and old rules can be directly used as $\mathcal{H}^{(t+1)}$, if it doesn’t exceed the capacity limit $H$, otherwise we need to discard some rules based on their classification accuracy. The rule weights $\theta^{(t+1)}$ are also updated. This process is repeatedly executed for several times to iteratively refine the rules, while the whole process is monitored on $\mathcal{S}_{\mathrm{val}}$ for early stopping.
  4. 4.
    Evaluation: After getting generated rules, we conduct a comprehensive evaluation on $\mathcal{S}_{\mathrm{te}}$ using two classes of inference strategies: direct prediction using the linear combiner, and injecting the learnt rules and their weights into an LLM for augmented reasoning, to evaluate the quality of rules with different inference strategies.
3.1 Rule Generation
Initialization and Candidate Generation. In the first iteration, we initialize the rule set as empty, $\mathcal{H}^{(0)}:=\emptyset$. We randomly sample $k$ examples from $\mathcal{S}_{\mathrm{tr}}$ and prompt an LLM to generate $h$ candidate rules (where $h\leq H$), denoted as $\mathcal{H}_{\mathrm{cand}}^{(1)}=\{{h}^{(1)}_{1},\ldots,{h}_{h}^{(1)}\}$.
Individual Rule Application. For any samples $x_{i}$ and candidate rule ${h}_{j}^{(1)}$, we prompt an LLM to obtain a local, ternary judgment
$z_{i,j}^{(1)}=\mathrm{LLM}\left(x_{i},{h}_{j}^{(i)}\right)\in\{-1,0,+1\},$
as $z_{i,j}$ means the evaluation results of applying the $j$-th rule on the $i$-th sample where +1+1 indicates a positive prediction, 1-1 a negative prediction and 0 signifies the rule is not applicable (“abstain”). Allowing abstention is crucial for explicitly modeling rule coverage, reducing misclassifications, and enabling sparse, robust combinations in the subsequent stage.
Coverage-based Filtering and Rule Set Update. For each candidate rule ${h}_{j}^{(1)}$, we compute its coverage on the training set as
$\mathrm{Cov}_{\mathrm{tr}}\left({h}_{j}^{(1)}\right)=\frac{1}{N_{\mathrm{tr}}}\sum_{i=1}^{N_{\mathrm{tr}}}\mathbb{I}\left[z_{i,j}^{(1)}\neq 0\right],$
and discard rules with coverage lower than a predefined threshold $\gamma$. After $\mathrm{Cov}_{\mathrm{tr}}$ based filtering, the valid rules generated here is denoted as $\mathcal{H}_{\mathrm{new}}^{(1)}$, and the rule set is updated as $\mathcal{H}^{(1)}=\mathcal{H}^{(0)}\cup\mathcal{H}_{\mathrm{new}}^{(1)}$.
3.2 Logistic Regression
Feature Construction. For any iteration $t$, given a rule set $\mathcal{H}^{(t)}=\{h_{1}^{(t)},\ldots,h^{(t)}_{m^{(t)}}\}$ where $m^{(t)}\leq H$, we define a mapping $\Phi^{(t)}$ from an input sample to a vector of rule judgments
$\Phi^{(t)}:\ \mathcal{X}\to\{-1,0,+1\}^{m^{(t)}},\qquad\Phi^{(t)}(x_{i})=\mathbf{z}_{i}^{(t)}=\begin{bmatrix}z_{i,1}^{(t)},\ldots,z_{i,m^{(t)}}^{(t)}\end{bmatrix}^{\top}$.
Weight Learning. The probability of a sample $x_{i}$ being classified as positive is modeled by a logistic regression function, which is defined as below
$p^{(t)}(x_{i};\theta^{(t)})=\sigma\!\big((\Phi^{(t)}(x_{i}))^{\top}\bm{\beta}^{(t)}+b^{(t)}\big),\qquad\sigma(u)=(1+e^{-u})^{-1}$
where $\sigma$ is the sigmoid function and $\theta^{(t)}=(\bm{\beta}^{(t)},b^{(t)})$ are the model parameters. We learn these parameters by minimizing the cross-entropy loss on $\mathcal{S}_{\mathrm{tr}}$ with Elastic Net (Zou & Hastie, 2005) regularization:
$\bm{\beta}^{(t)},b^{(t)}=\underset{{\bm{\beta}^{(t)},b^{(t)}}}{\arg\min}\ \frac{1}{|\mathcal{S}_{\mathrm{tr}}|}\!\sum_{x_{i}\in\mathcal{S}_{\mathrm{tr}}}\mathcal{L}(y_{i},p^{(t)}(x_{i};\theta^{(t)}))+\lambda\!\left(\alpha\|\bm{\beta}^{(t)}\|_{1}+\tfrac{1-\alpha}{2}\|\bm{\beta}^{(t)}\|_{2}^{2}\right),$
where $\mathcal{L}$ is the binary cross-entropy loss. The $L_{1}$ penalty encourages sparsity (rule selection), while the $L_{2}$ penalty enhances robustness. The regularization hyperparameters $(\lambda,\alpha)$ are selected via stratified K-fold cross-validation on $\mathcal{S}_{\mathrm{val}}$. The final parameters $\hat{\theta}^{(t)}=(\hat{\bm{\beta}}^{(t)},\hat{b}^{(t)})$ are obtained by refitting the model on the entire $\mathcal{S}_{\mathrm{tr}}$ with the selected hyperparameters.
Label Prediction. After getting rule weights, the prediction for a sample $x_{i}$ can be calculated as
$\hat{y}_{i}^{(t)}=\mathbb{I}\big[p^{(t)}(x_{i};\hat{\theta}^{(t)})\geq\tau\big]\in\{0,1\},\qquad\text{with }\tau=0.5.$
3.3 Iterative Refinement
In the first iteration, $\mathcal{H}^{(1)}$ and $\hat{\theta}^{(1)}$ are obtained on randomly selected samples. In the subsequent iterations, given $\mathcal{H}^{(t)}$ and $\hat{\theta}^{(t)}$ with $t\geq 1$, to generate $\mathcal{H}^{(t+1)}$ and $\hat{\theta}^{(t+1)}$, instead of relying on randomly selected samples, we can enhance the rule generation by targeting hard examples.
1. Hard Example Selection. For each example $x_{i}\in\mathcal{S}_{\mathrm{tr}}$, we compute its predicted probability $\hat{p}_{i}^{(t)}=p^{(t)}(x_{i};\hat{\theta}^{(t)})$ and define its prediction error as $d_{i}=\|\hat{p}_{i}^{(t)}-y_{i}\|$. We identify the top-$k$ samples with the highest error $I_{\mathrm{hard}}^{(t)}=\mathrm{TopK}\left(\{(i,d_{i})\}^{\|\mathcal{S}_{\mathrm{tr}}\|}_{i=1},k\right)$.
2. New Rule Generation. The selected hard examples $I^{(t)}_{\mathrm{hard}}$ are presented to the LLM along with the current rule set $\mathcal{H}^{(t)}$. The LLM is prompted to reflect on the errors and either revise existing rules or generate new ones that cover novel perspectives. The $h$ newly generated rules are filtered for coverage, resulting in the set $\mathcal{H}_{\mathrm{new}}^{(t+1)}$.
3. Rule Set Update. The new rules are merged with the existing set: $\mathcal{H}_{\mathrm{tmp}}^{(t+1)}=\mathcal{H}^{(t)}\cup\mathcal{H}_{\mathrm{new}}^{(t+1)}$. If the size of this temporary set does not exceed the capacity $H$, it can be directly used as $\mathcal{H}^{(t+1)}$, otherwise we have to prune the rules by ranking them based on their individual accuracy on the validation set, retaining only the top $H$ rules to form $\mathcal{H}^{(t+1)}$.
4. Parameter Update and Termination. With the updated rule set $\mathcal{H}^{(t+1)}$, we calculate $\hat{\theta}^{(t+1)}$ in the same way as discussed in Section 3.2, after which we can monitor the overall performance on $\mathcal{S}_{\mathrm{val}}$. Rules will be refined iteratively, with each iteration trying to improve the performance on hard examples of previous version of rule sets. The process terminates if the overall performance on $\mathcal{S}_{\mathrm{val}}$ fails to improve by a margin $\delta$ for $p$ consecutive iterations or if the maximum number of iterations $R_{\mathrm{max}}$ is reached. The final model $(\mathcal{H}^{\star},\hat{\theta}^{\star})$ corresponds to the checkpoint with the best validation performance.
3.4 Evaluation
We systematically compare four inference strategies on the test set $\mathcal{S}_{\mathrm{te}}$, categorized into two main approaches: direct inference using the linear combiner and LLM-augmented inference with varying levels of information.
(E1) Linear-only. The final prediction is made directly by the logistic regression model, with $\hat{y}_{i}^{\mathrm{final}}=\mathbb{I}\big[p(x_{i};\hat{\theta}^{\star})\geq 0.5]$, where $p(\cdot;\hat{\theta}^{\star})$ is the learnt probabilistic combiner. This serves as a classic baseline for weighted rule sets.
(E2) LLM + Rules. The LLM is provided with only the texts of the rules in $\mathcal{H}^{\star}$ and the input $x_{i}$, and is tasked with making a prediction where $\hat{y}_{i}^{\mathrm{final}}=\mathrm{LLM}(x_{i},\mathcal{H}^{\star})$. This evaluates the LLM’s intrinsic ability to aggregate rule-based evidence, mirroring common practices in existing work.
(E3) LLM + Rules + Weights. In addition to the rule texts, the LLM is also provided with the learnt weights $\hat{\theta}^{\star}$ and the final prediction becomes $\hat{y}_{i}^{\mathrm{final}}=\mathrm{LLM}(x_{i},\mathcal{H}^{\star},\hat{\theta}^{\star})$. This tests whether the LLM can leverage explicit probabilistic signals to guide its reasoning, reflecting a “probabilistic internalization” of rule importance.
(E4) LLM + Rules + Weights + Linear Prediction. The LLM receives all information from generated rules and corresponding weights, plus the prediction $\hat{y}_{i}$ from the linear-only model as a reference, so the prediction becomes $\hat{y}_{i}^{\mathrm{final}}=\mathrm{LLM}(x_{i},\mathcal{H}^{\star},\hat{\theta}^{\star},\hat{y}_{i})$. This strategy assesses whether a potentially imperfect but calibrated rule-based system can provide a valuable external signal to enhance the LLM’s reasoning in complex scenarios where linear assumption may not fully hold.
4 Experiment Setup
4.1 Dataset
To evaluate our method, we select six real-world tasks from the HypoBench(Liu et al., 2025) benchmark, including Deception Detection (Review), Persuasive Argument Prediction (Dreddit), News Headline Engagement (Headlines), Paper Citations (Citations), AI-generated Content Detection (LLM Detect) and Retweets (Retweets). Each of the task can be viewed as a binary classification task, with text information as inputs and binary classification results as output (e.g. given a review, we need to evaluate whether it is deceptive or not). Details of each dataset are in Appendix A.3.
4.2 Baseline
We implemented and compared RLIE against several representative and competitive methods within a unified framework to comprehensively evaluate its performance and robustness. (1) Zero-shot Inference: The LLM directly predicts the output for a given sample based only on the task description, without access to any examples or rules. (2) Zero-shot Generation: The LLM generates a set of candidate rules based solely on the task description. The single best-performing rule, as determined on the validation set, is then used for inference on the test set. (3) IO Refinement (Qiu et al., 2023): This method follows a “propose-select-refine” loop. In each round, the LLM generates a batch of candidate hypotheses, the best one is selected based on its performance on a validation or error set, and it is then used as a seed for the next round of refinement. This process is repeated for a fixed number of iterations. (4) HypoGeniC (Zhou et al., 2024): This approach maintains and dynamically expands a library of hypotheses. For each sample, the top-$k$ hypotheses are selected to make a prediction. Mismatched samples are added to an error set, which triggers the generation of new hypotheses to improve coverage. The hypothesis library is continuously updated based on a reward signal.
4.3 Experimental Details
In our experiments, each dataset was partitioned into fixed-size training, validation, and test sets of 200, 200, and 300 samples, respectively. All methods used the same data splits and a fixed random seed to ensure reproducibility. We report Accuracy and Macro-F1 score. Each experiment was repeated at least three times, and we report the mean and standard deviation of the results. All experiments involving LLMs utilized gpt-4o-mini with the temperature set to $1\times 10^{-5}$ to ensure deterministic outputs. Prompt details can be found in Appendix C. As for hyperparameters, the capacity of the rule set was set to $H=10$. In each iteration, we sampled $k=20$ hard examples and generated $h=5$ new rules. The minimum coverage threshold for rules was set to $\gamma=0.2$. The implementations of IO Refinement and HypoGeniC followed the default hyperparameter settings of the original papers. Further details are provided in Appendix A.
5 Results
This section presents empirical evaluation of our framework. First, we compare the overall performance of RLIE against competitive baselines. Second, we conduct an ablation study on different inference strategies to identify the most effective way to utilize the learned probabilistic rules.
Table 1: Overall performance of different frameworks evaluated in Accuracy and F1 Score ($mean\pm std$, $\times 10^{-2}$). The best-performing results are shown in bold, while the second best results are highlighted in underline. The performance of RLIE is evaluated with the Linear Only strategy.
Method Reviews Dreaddit Headlines Citations LLM Detect Retweets
Accuracy
Zero-shot Inference 55.11±0.18 62.00±0.88 63.22±2.52 56.25±0.00 56.11±0.38 65.56±2.34
Zero-shot Generation 63.89±5.50 64.33±2.19 53.67±1.20 58.33±3.61 52.44±1.07 57.67±2.40
IO Refinement 64.11±5.52 74.89±3.47 59.67±3.51 48.96±1.80 60.33±17.32 62.44±0.69
HypoGeniC 58.00±6.35 65.67±0.00 60.78±3.27 47.92±7.22 54.56±4.43 65.56±2.12
RLIE (Ours) 71.00±0.67 69.11±1.17 65.56±1.50 59.38±3.12 58.67±2.65 65.78±1.71
F1 Score
Zero-shot Inference 50.36±2.30 55.67±2.30 62.70±2.53 36.00±0.00 46.77±0.68 64.69±2.20
Zero-shot Generation 63.41±6.09 59.17±2.84 53.10±0.66 42.56±7.04 39.68±1.58 57.64±2.42
IO Refinement 62.03±8.45 73.98±3.59 59.21±3.23 45.66±2.19 54.00±22.42 62.30±0.66
HypoGeniC 56.82±6.29 60.57±0.61 59.56±3.82 41.96±3.40 44.17±8.75 65.49±2.17
RLIE (Ours) 70.81±0.59 65.94±1.25 65.38±1.44 49.08±2.84 52.10±4.51 65.45±1.41
5.1 Main Results
The empirical results, summarized in Table 1, demonstrate the effectiveness and robustness of RLIE compared to several strong baselines across six real-world datasets. For this comparison, RLIE’s performance is evaluated using its direct inference strategy (Linear-only), where the learned logistic regression model acts as the classifier. While some baseline methods may attain the top performance on one or two individual datasets, RLIE consistently ranks within the top two in terms of both Accuracy and F1 Score. This demonstrates that the RLIE framework can be generalized to diverse data distributions and is effective at capturing underlying data patterns. As shown in the results table, some methods exhibit high variance and instability despite their strong performance on certain tasks (e.g., IO Refinement). In contrast, our method achieves high performance while maintaining stability, underscoring its robustness.
A horizontal comparison of the different methods reveals several interesting insights. It is noteworthy that Zero-shot Inference outperforms many of the more complex baselines in several scenarios. This may be attributed to the extensive knowledge acquired by the LLM during pre-training, which is sufficient for solving certain tasks. However, this internal inference process is inherently unexplainable, and it is difficult to rule out the influence of potential data leakage from the pre-training corpus. This observation highlights both the challenge and the necessity of using rule-based learning to achieve interpretable reasoning. HypoGeniC learns a set of rules independently and fails to account for the interactions between them, resulting in suboptimal performance. By iteratively refining a single rule, IO Refinement achieves respectable results, which validates the value of providing annotated examples and employing an iterative optimization process. It is worth mentioning that in some cases, IO Refinement outperforms RLIE. This could be because the strategy of generating only a single rule forces it to be more generalizable, allowing it to better handle unseen scenarios. However, this single-rule approach also limits the method’s expressiveness, leading to poor performance in cases that require a more complex rule set. Our RLIE framework strikes a balance: by continuously learning from hard examples, it enhances the generalizability of the learned rule set, while the strategy of maintaining multiple rules significantly boosts its expressiveness. This allows our method to achieve excellent and robust performance with low variance across diverse datasets.
Table 2: Performance of different inference strategies valuated in Accuracy and F1 Score ($mean\pm std$, $\times 10^{-2}$). The best-performing results are shown in bold, while the second best results are highlighted in underline.
Inference Strategy Reviews Dreaddit Headlines Citations LLM Detect Retweets
Accuracy
Regression Only 71.00±0.67 69.11±1.17 65.56±1.50 59.38±3.12 58.67±2.65 65.78±1.71
LLM Evaluation 65.22±2.67 64.56±0.51 61.33±3.06 60.42±1.80 55.22±2.12 61.00±3.06
LLM + Linear Regression 64.67±2.33 67.56±2.50 61.89±3.02 60.42±1.80 54.11±2.22 62.22±3.56
LLM + Linear + Label 64.67±2.60 69.00±1.33 65.33±1.86 59.38±3.12 56.67±3.51 60.22±3.56
F1 Score
Regression Only 70.81±0.59 65.94±1.25 65.38±1.44 49.08±2.84 52.10±4.51 65.45±1.41
LLM Evaluation 64.52±3.19 59.57±0.94 60.88±3.18 48.62±1.19 46.82±5.32 60.33±3.13
LLM + Linear Regression 63.87±3.05 63.73±3.72 61.71±2.93 47.11±3.38 44.76±4.62 59.68±4.45
LLM + Linear + Label 63.80±2.87 65.79±1.51 65.16±1.78 49.08±2.84 48.70±6.05 57.38±4.84
5.2 How to Best Utilize Rules?
To determine the optimal strategy for leveraging the learned probabilistic rules, we evaluated the four distinct inference methods detailed in Section 3.4. One might hypothesize that an LLM, with its vast pre-trained knowledge, could synthesize the rule set, their learned weights, and even a reference prediction to produce superior judgments. However, the empirical results, presented in Table 2, are counterintuitive.
The most striking finding is that the simplest strategy, Linear-only, where the logistic regression model makes the final prediction, achieves the best performance on nearly all datasets. It outperforms all three LLM-based inference strategies in most cases, particularly in terms of F1 score where it is the top performer across all six tasks. This highlights the high quality of the rule set curated by RLIE and the effectiveness of the probabilistic combiner.
Injecting more information into the LLM does not yield consistent benefits. Providing the LLM with rule weights (LLM + Rules + Weights) over just the rule texts does not lead to stable performance gains, suggesting that the LLM struggles to effectively internalize and apply probabilistic information to refine its reasoning. Most surprisingly, providing the LLM with the linear model’s own (often correct) prediction as a reference (LLM + Rules + Weights + Linear Prediction) frequently results in performance degradation compared to simply using the linear model’s prediction directly. This suggests that the LLM, when tasked with integrating qualitative probabilities and multiple form evidences into reasoning, can be led astray, potentially overwriting a correct judgment with an incorrect one. This finding aligns with existing research indicating that LLMs can be inconsistent when required to strictly adhere to complex, explicit instructions, showing the deficiency of LLMs to perform fine controlled inference.
6 Discussion
Our hierarchical evaluation reveals a central finding: achieving stable, globally consistent reasoning by prompting an LLM with a set of explicitly weighted rules is non-trivial. The simplest “Linear-only” strategy, which uses the learned logistic regression model as the final classifier, performed best on nearly all datasets and achieved the highest F1 score across the board. In contrast, injecting the rule texts, their weights, and even the linear model’s reference prediction back into the LLM often failed to yield stable improvements and, in some cases, degraded performance. This result aligns with the observation that LLMs excel at semantic generation and interpretation but are less reliable at fine-grained, controlled probabilistic integration. This suggests that a more robust paradigm involves a clear division of labor: leveraging LLMs for local, semantic tasks such as judging individual rules and generating new candidates, while entrusting a transparent and calibratable probabilistic combiner—in our work, a logistic regression model with elastic net regularization—with the global tasks of weighting and selection.
From a broader perspective, using natural language rules as the fundamental unit of reasoning bypasses the rigid dependency on predefined predicate spaces found in traditional methods, offering greater expressiveness for open-ended semantics and unstructured data. At the same time, this approach avoids overwhelming the LLM’s context with complex probabilistic details from an entire rule set, thereby preserving its ability to make stable judgments at the single-rule level. Unlike frameworks that model interactions at the predicate level (e.g., Markov Logic Networks), we advocate for enhancing the capabilities of the global combiner while keeping the local interface—the LLM’s ternary judgments and coverage information—fixed. This linear log-odds combiner can be naturally extended to more sophisticated models, such as (i) interpretable additive models (e.g., GAMs) to capture rule interactions; (ii) message passing on a factor graph to globally harmonize rule confidences; (iii) Bayesian logistic regression with hierarchical priors (e.g., ARD, Laplace) for joint uncertainty and sparsity modeling; or (iv) techniques like Platt scaling or isotonic regression to explicitly calibrate the LLM’s ”abstain” outputs. These extensions are fully compatible with the current RLIE architecture. Overall, we propose a replicable engineering principle: LLMs are responsible for semantic generation and local judgment, while traditional probabilistic learning methods handle global aggregation and uncertainty management. On this foundation, stronger graphical or additive structures and more advanced inference algorithms can be progressively introduced to build more flexible and powerful, yet still interpretable and calibratable, neuro-symbolic systems.
7 Conclusion
In this paper, we addressed the critical challenge of learning robust and interpretable rule sets from data. While Large Language Models (LLMs) have created new possibilities for generating rules in natural language, existing methods have struggled with the principled aggregation and calibration of these rules. We introduced RLIE, a unified framework that synergizes the rule-generation capabilities of LLMs with the global optimization of probabilistic models. By integrating rule generation, regularized logistic regression, and error-driven iterative refinement, RLIE produces a compact and calibrated set of weighted, natural language rules. Our experiments demonstrate that this hybrid approach is highly effective. However, directly using LLMs to apply weighted rules degrades the performance of downstream tasks, which further unveils the limitations of applying LLMs in rule based reasoning. Our work underscores the value of a neuro-symbolic division of labor, where LLMs handle local semantic tasks and classical models manage global probabilistic reasoning. The paper investigates the potentials and limitations of using LLMs for inductive reasoning tasks, proposing a unified framework which integrates LLMs with classic probabilistic rule combination methods, paving the way for more building reliable AI and contribute to knowledge discovery.
8 Ethic Statement
This research adheres to the ICLR Code of Ethics. The datasets used in our experiments are derived from the HypoBench benchmark, which is composed of publicly available data from prior academic research. To the best of our knowledge, personally identifiable information has been anonymized in these datasets. We acknowledge that the rules generated by our method reflect the patterns and potential societal biases present in the underlying training data. While a primary goal of our work is to enhance interpretability, this also means that any learned biases will be transparent. We caution that the application of such rule-based systems, particularly in sensitive domains like mental health analysis or persuasion, should be accompanied by careful human oversight to mitigate the risk of misuse or the perpetuation of harmful stereotypes. Our work is intended to contribute to the development of more transparent and auditable AI systems.
9 Reproducibility Statement
We are committed to ensuring the reproducibility of our research. The complete source code for the RLIE framework, along with the scripts for running all experiments, will be made publicly available upon publication. All datasets used in this study are from the public HypoBench benchmark, and we will provide the exact data splits to facilitate direct comparison. A detailed description of our experimental setup, including model specifications, baseline implementations, and primary hyperparameters, is provided in Section 4. A comprehensive list of all prompts used to interact with the LLM for rule generation, judgment, and inference is included in the Appendix C. The code will be released once the paper is accepted.
Appendix A More Details
In this section, we provide additional information regarding the baseline methods used for comparison and the specific validation procedures employed in our experiments.
A.1 Baselines
To contextualize the performance of RLIE, we compare it against two state-of-the-art LLM-based rule learning frameworks:
  • IO Refinement: Proposed by Qiu et al. (2023), this method follows a ”propose-select-refine” loop. It begins by generating candidate rules and iteratively refines a single best hypothesis based on feedback from its performance on a set of examples.
  • HypoGeniC: Introduced by Zhou et al. (2024), this framework maintains a dynamic library of hypotheses. It uses a reward mechanism inspired by multi-armed bandits to balance the exploration of new rules and the exploitation of well-performing ones, updating its rule set based on performance on training instances.
A.2 Validation
Our experimental setup follows a standardized procedure, but certain details regarding the use of the validation set for baselines and the handling of smaller datasets warrant clarification.
For datasets where the total number of available samples was less than the specified 200 for training and 200 for validation, we utilized all available data for the respective splits.
The baseline methods handle the validation set differently according to their original designs. Since the algorithms for IO Refinement and Zero-shot Generation do not inherently use a validation set during their rule generation phase, we implemented a final selection step for a fair comparison: after these methods completed their final iteration, we selected the single rule that achieved the highest performance on our designated validation set to be used for testing. In contrast, the HypoGeniC algorithm does not involve a validation set in its update loop; its mechanism for updating and pruning the hypothesis library relies on rewards calculated from performance on the training batches. Therefore, for HypoGeniC, the validation set was not used during its training and refinement process.
A.3 Dataset
Deception Detection (Reviews). The objective is to distinguish between genuine and deceptive hotel reviews. Deceptive reviews, often written by paid individuals, may appear plausible but typically lack specific details. The primary challenge lies in identifying subtle linguistic cues, such as exaggerated sentiment or vague descriptions, which are difficult for even human evaluators to discern reliably.
Persuasive Argument Prediction (Dreddit). Given a pair of argumentative texts, the model must predict which is more persuasive. This task emphasizes linguistic features correlated with persuasiveness, such as argument structure, rhetorical devices, and emotional tone. It is particularly challenging as persuasiveness often depends on complex context and psychological factors.
News Headline Engagement (Headlines). Given two headlines for the same news event, the model must predict which one is more likely to attract clicks. The task examines how linguistic choices in news writing—such as novelty, specificity, or emotional framing—influence reader behavior. The often minimal difference between headlines makes this task highly challenging.
Paper Citations (Citations). This task involves predicting whether an academic paper will achieve a high citation count based on its title and abstract. It focuses on features related to academic impact, such as the generality of the research problem, the novelty of the contributions, or the timeliness of the research topic. Unlike social media tasks, this task emphasizes long-term value signals in academic writing.
AI-generated Content Detection (LLM Detect). This task requires determining whether a given text was written by a human or generated by an LLM. The dataset contains human- and machine-written texts based on the same prompts, evaluating a model’s ability to capture stylistic, structural, or semantic differences indicative of AI-generated content and formulate them as interpretable rules.
Retweets (Retweets). The input is a pair of tweets, and the task is to predict which one will be retweeted more. Influential factors include a tweet’s conciseness, emotional intensity, and the mention of specific individuals or organizations. Due to the highly stochastic nature of social media propagation, this task places stringent demands on the explanatory power and stability of the learned rules.
Appendix B LLM Usage
Large Language Models (LLMs) were used as an assistive tool in preparing this manuscript. The core intellectual contributions, including the research ideas, experimental design, and analysis, are entirely the work of the human authors. The LLM’s role was limited to language polishing and coding assistance. The authors have reviewed and edited all LLM-generated content and take full responsibility for the final manuscript.
Appendix C Prompts
Here we are providing the prompts we are using to conduct all the experiments. For the clarity of illustration, here we are only showing prompts for the Retweet task, while all the prompts can be access after this paper is accepted.
multi_content: |
The first tweet: \${first_tweet}
The second tweet: \${second_tweet}
Final answer: The \${label} tweet got more retweets.
Figure 2: Prompt for providing observations.
system: |-
You are a social media expert. You are an expert at determining which tweet will be retweeted more.
Given a set of observations, you want to generation hypotheses that will help predict which tweet out of a pair of tweets is more likely to be retweeted.
Please note that the paired tweets are about the same content and are posted by the same user, so you should focus on the wording difference between the two tweets in each pair.
Please propose \${num_hypotheses} possible hypotheses.
Please generate them in the format of:
1. [hypothesis]
2. [hypothesis]
\${num_hypotheses}. [hypothesis].
Please make the hypotheses general enough to be applicable to new observations.
user: |-
We made some observations:
\${observations}
Generate hypotheses that are useful for predicting which tweet out of a pair of tweets is more likely to be retweeted.
Please note that the paired tweets are about the same content and are posted by the same user, so you should focus on the wording difference between the two tweets in each pair.
Please propose \${num_hypotheses} possible hypotheses.
Please generate them in the format of:
1. [hypothesis]
2. [hypothesis]
\${num_hypotheses}. [hypothesis].
Proposed hypotheses:
Figure 3: Prompt for the first iteration of rule generation.
system: |-
You are a social media expert focused on maximizing retweet engagement.
Given misclassified tweet pairs and prior hypotheses, your goal is to rethink and propose \${num_hypotheses} new, more accurate hypotheses about which tweet in a pair will earn more retweets.
Please note that the paired tweets share the same content and author, so concentrate on wording differences, framing, and presentation.
Generate the hypotheses in the format of:
1. [hypothesis]
2. [hypothesis]
\${num_hypotheses}. [hypothesis].
Please make the hypotheses general enough to be applicable to new observations.
user: |-
We have tweet pairs that previous hypotheses predicted incorrectly:
\${observations}
Here are some of the prior hypotheses for reference:
\${hypotheses_text}
Please generate new hypotheses that better capture which tweet in each pair will get more retweets.
You may refine the previous hypotheses (like tightening conditions, adding exceptions, or rephrasing), or introduce new hypotheses to cover new, distinct angles when prior ones are insufficient or misaligned.
Propose \${num_hypotheses} possible hypotheses.
Generate them in the format of 1. [hypothesis], 2. [hypothesis], ... \${num_hypotheses}. [hypothesis].
Proposed hypotheses:
Figure 4: Prompt for iterative refinement.
system: |-
You are a social media expert.
Given a pair of tweets, you are asked to predict which tweet will be retweeted more.
Please note that the paired tweets are about the same content and are posted by the same user, so you should focus on the wording difference between the two tweets.
From past experiences, you learned a pattern.
Now, at each time, you should apply a learned pattern to a pair of tweets and determine which one will get more retweets.
Give an answer. Respond with exactly one of: first, second, or not applicable.
Give your final answer in the format of {Final answer: first} or {Final answer: second}.
user: |-
Pattern: \${hypothesis}
The first tweet: \${first_tweet}
The second tweet: \${second_tweet}
Given the pattern you learned above, predict which one of the two tweets will get more retweets.
Think step by step.
First step: Consider if the pattern can be applied to analyze the textual difference between the two tweets.
Third step: Based on the pattern, which tweet is more likely to get more retweets? If it does not apply, say so explicitly.
Final step: Give your final answer in the format of {Final answer: first}, {Final answer: second} or {Final answer: not applicable}.
Final answer:
Figure 5: Prompt for single hypothesis inference.
system: |-
You are a social media expert.
Given a pair of tweets, you are asked to determine which will get more retweets.
From past experiences, you learned some patterns.
You need to determine whether each of the patterns holds for the current pair of tweets, and also predict which tweet will get more retweets.
Give your final answer in the format of {Final answer: first} or {Final answer: second}.
user: |-
Our learned patterns: \${hypotheses}
The first tweet: \${first_tweet}
The second tweet: \${second_tweet}
Given the patterns you learned above, predict which one will get more retweets.
Think step by step.
First step: Think about which patterns can be applied to these tweets.
Second step: Based on the applicable patterns, which tweet is likely to get more retweets?
Give your final answer in the format of {Final answer: first} or {Final answer: second}.
Figure 6: Prompt for “LLM + Rules” style inference.
system: |-
You are a social media expert.
Given a pair of tweets, you are asked to determine which will get more retweets.
We trained a linear regression model on the training set to obtain a collection of weighted patterns plus a bias term. The learned weight reflects how strongly the pattern contributes to predicting "first" vs "second".
Review the weighted hypotheses, consider how the bias interacts with them, and use the regression models suggested label only as a reference.
Give your final answer in the format of {Final answer: first} or {Final answer: second}.
user: |-
Our learned weighted patterns (the weights magnitude reflects the patterns importance):
\${weighted_hypotheses}
Bias term (a constant offset added regardless of pattern activations; a positive bias means the model is overall more inclined toward \${pos_label}, while a negative bias leans toward \${neg_label}):
\${bias}
The first tweet: \${first_tweet}
The second tweet: \${second_tweet}
Given the patterns you learned above, predict which one will get more retweets.
Think step by step.
First step: Think about which patterns can be applied to these tweets.
Second step: Decide which tweet is likely to get more retweets, you can used the weighted patterns and bias as reference.
Final step: give your final answer in the format of {Final answer: first} or {Final answer: second}.
Figure 7: Prompt for “LLM + Rules + Weights” style inference.
system: |-
You are a social media expert.
Given a pair of tweets, you are asked to determine which will get more retweets.
We trained a linear regression model on the training set to obtain a collection of weighted patterns plus a bias term. The learned weight reflects how strongly the pattern contributes to predicting "first" vs "second". The regression model also outputs a referenced label for the review, which should be treated as a suggestion.
Review the weighted hypotheses, consider how the bias interacts with them, and use the regression models suggested label only as a reference.
Give your final answer in the format of {Final answer: first} or {Final answer: second}.
user: |-
Our learned weighted patterns (the weights magnitude reflects the patterns importance):
\${weighted_hypotheses}
Bias term (a constant offset added regardless of pattern activations; a positive bias means the model is overall more inclined toward \${pos_label}, while a negative bias leans toward \${neg_label}):
\${bias}
The first tweet: \${first_tweet}
The second tweet: \${second_tweet}
We have used the regression model to get a referenced label: \${model_prediction}
Given the patterns you learned above, predict which one will get more retweets.
Think step by step.
First step: Think about which patterns can be applied to these tweets.
Second step: Decide which tweet is likely to get more retweets, you can used the weighted patterns, bias, predicted label as reference.
Final step: give your final answer in the format of {Final answer: first} or {Final answer: second}.
Figure 8: Prompt for “LLM + Rules + Weights + Linear Prediction” style inference.

### v2

@@
 Abstract
-Nowadays, Large Language Models (LLMs) are able to propose rules in natural language, overcoming constrains of a predefined predicate space inherent in traditional rule learning.
-However, existing methods using LLMs often overlook the combination effects of rules, and the potential of coupling LLMs with probabilistic rule learning to ensure robust inference is not fully explored.
-To address this gap, we introduce RLIE, a unified framework that integrates LLMs with probabilistic modeling to learn a set of probabilistic rules.
-The RLIE framework comprises four stages: (1) Rule generation, where a LLM proposes and filters candidate rules; (2) Logistic regression, which learns the probabilistic weights of the rules for global selection and calibration; (3) Iterative refinement, which continuously optimizes the rule set based on prediction errors; and (4) Evaluation, which compares the performance of the weighted rule set as a direct classifier against various methods of injecting the rules into an LLM.
-Generated rules are the evaluated with different inference strategies on multiple real-world datasets.
-While applying rules directly with corresponding weights brings us superior performance, prompting LLMs with rules, weights and classification results from the logistic model will surprising degrade the performance.
-This result aligns with the observation that LLMs excel at semantic generation and interpretation but are less reliable at fine-grained, controlled probabilistic integration.
-Our work investigates the potentials and limitations of using LLMs for inductive reasoning tasks, proposing a unified framework which integrates LLMs with classic probabilistic rule combination methods, paving the way for more reliable neuro-symbolic reasoning systems.
+Large Language Models (LLMs) can propose natural-language rules, circumventing the reliance on a predefined predicate space in traditional rule learning.
+However, existing LLM-based methods often neglect the global interactions among rules, and the potential of using fine-grained rule importance scores to calibrate neuro-symbolic reasoning remains underexplored.
+To address this gap, we introduce RLIE, a framework that integrates LLMs with probabilistic modeling to learn weighted rule sets in four stages: (1) Rule generation: proposing and filtering candidate rules via LLMs; (2) Logistic regression: learning sparse, calibrated weights for global rule selection; (3) Iterative refinement: revising the rule set with error-driven hard examples; and (4) Evaluation: validating the learned system via comparative inference paradigms.
+Across multiple real-world datasets and LLM backbones, our learned weighted rules achieve superior stability and accuracy, whereas rule-injection prompting yields mixed results and often degrades performance.
+These results suggest LLMs excel at semantic rule discovery but are less reliable at controlled probabilistic aggregation.
+Our findings highlight both the promise and the limits of LLMs for inductive reasoning, motivating a principled integration with classic probabilistic rule combination for reliable neuro-symbolic reasoning.
+Machine Learning, ICML
 1 Introduction
-In data-driven applications and scientific discovery, the goal is not merely to predict outcomes, but to construct a set of verifiable, reusable, and composable theories(Zhou et al., 2024; Yang et al., 2024a; Minh et al., 2022).
-These theories can enable explainable, auditable decisions while driving the discovery of new knowledge and underlying structures(Yang et al., 2023; 2024b).
-These theories can be expressed in formal, structural statements(Cohen et al., 1995; Cropper & Morel, 2021) or natural language hypotheses(Zhou et al., 2024), and they share a common characteristic: they are declarative, testable, and self-contained discriminative patterns that yield predictions verifiable by external evidence [^1] In this paper, we do not distinguish between the terms “rule” and “hypothesis”, and will use “rule” throughout the text for consistency..
-The challenge of how to learn and integrate a set of such collaborative patterns to both uncover regularities and advance knowledge accumulation has long been a central problem in machine learning and artificial intelligence(Bazgir et al., 2025).
-Take the binary classification task of spam detection as an example.
-A set of rules might include (1) $\mathrm{HasToken}(\text{`win prize'},x)\land\mathrm{HasAttachment}(x)$; (2) $\mathrm{DomainInBlacklist}(x)\land\mathrm{HasShortLink}(x)$; and (3) $\mathrm{TitleAllCaps}(x)\land\mathrm{HasToken}(\text{`urgent'},x)$, all implying the head “is spam”.
-These rules cover different facets of the problem, and must work in concert.
-A classic approach to use the rules to do reasoning is deterministic aggregation, where a positive classification is made, if any rule is satisfied(Fürnkranz & Kliegr, 2015).
-A probabilistic approach, in contrast, treats each rule’s classification result as a binary indicator ($r_{j}\in\{0,1\}$) and employs a statistical model, such as logistic regression, to learn a global set of weights for combining these signals into a unified prediction(Ruczinski et al., 2003; Friedman & Popescu, 2008).
-While the well-defined structure of rules offers significant advantages, traditional rule learning methods are constrained by a predefined predicate space, which limits their ability to handle unstructured data and open-ended semantics, thereby restricting their applicability in broader scenarios(Cerna & Cropper, 2024; Hocquette et al., 2024).
-The recent advent of Large Language Models (LLMs) presents a new opportunity for rule learning.
-Given a few examples and task context, LLMs can directly generate rules in natural language(Ellis, 2023; Singh et al., 2022).
-Their linguistic richness and contextual alignment facilitate the articulation of fine-grained conditions, relations and exceptions, covering patterns that are difficult to capture with manually defined predicates.
-However, existing methods often focus on the iterative optimization of a single rule(Qiu et al., 2023) or the independent generation and simple aggregation of multiple rules(Yang et al., 2023; Zhou et al., 2024; Yang et al., 2024b), failing to jointly learn and compress the rules as a cohesive set.
-Concurrently, probabilistic approaches in the context of LLMs remains largely unexplored, as LLMs’ ability to robustly learn the rule weights from data and apply them to perform reliable probabilistic reasoning during inference is not systematically examined yet.
-This gap motivates our pursuit of a framework that combines the generative and expressive capabilities of LLMs with the probabilistic composition and calibration of traditional methods to produce a flexible, robust and reusable rule set.
-To achieve this synergy, we propose RLIE, a unified framework for LLM-based rule learning that integrates Rule generation, Logistic regression, Iterative refinement and Evaluation.
-First, in the rule generation stage, an LLM produces a set of independent, individually assessable natural language rules based on a few examples and the task context.
-Second, for probabilistic combination, a regularized logistic regression model is employed to perform global weighting and selection on the rules, yielding a compact, interpretable rule set with importance weights.
-Third, in the iterative refinement stage, challenging examples are identified based on the prediction errors.
-These examples, along with the current rule set, are fed back to the LLM to prompt reflective generation of improved rules until convergence.
-Finally, a evaluation of two primary inference strategies is conducted: (1) direct inference using the weighted model, and (2) injecting the rule information into an LLM.
-For the latter, we systematically examine three levels of information: (i) providing only the rule, (ii) providing the rules with their weights, and (iii) providing the rules, weights and the linear model’s prediction as a reference.
-The layered design allows for a systematic evaluation of quality of rules and analysis of the trade-offs in accuracy and robustness for different ways of using rules.
-Using our framework, we can get high quality rules in natural languages, with optimal ways of utilizing them to perform rule based reasoning.
-The main contribution of this paper are as follows:
-  1. 1.
-    We propose RLIE, a unified paradigm for LLM-based probabilistic rule learning.
-It combines the natural language rule generation of LLMs with the global weighting and selection of regularized logistic regression to produce a self-consistent, collaborative and auditable rule set.
-  2. 2.
-    We design a hierarchical set of inference strategies to systematically compare direct inference with the linear combiner against three methods of injecting information into an LLM: rules-only, rules+weights and rules+weights+linear prediction reference.
-We provide empirical analysis and practical guidance regrading their respective trade-offs in terms of accuracy, robustness and calibration.
-  3. 3.
-    On multiple real-world datasets, RLIE achieves superior over all performance compared to a range of LLM-based methods.
-The learnt rule sets are more compact and semantically clearer, prompting knowledge discovery and human-AI consensus.
+In many data-driven applications and scientific inquiry, the goal is increasingly shifting from pure prediction to constructing verifiable, reusable, and composable theories (Zhou et al., 2024; Yang et al., 2024a; Minh et al., 2022).
+Such theories enable explainable and auditable decisions and, in scientific contexts, can serve as building blocks for discovering underlying structures and accumulating knowledge (Yang et al., 2023, 2024b).
+These theories may be expressed as formal, structural statements (Cohen and others, 1995; Cropper and Morel, 2021) or as natural-language hypotheses [^1] In this paper, we use “rule” and “hypothesis” interchangeably, and refer to both as “rule”.
+(Zhou et al., 2024); regardless of form, they are declarative and testable units whose predictions can be verified by external evidence.
+A longstanding challenge is how to learn a collection of such testable hypotheses and integrate them into a coherent rule set that captures regularities, supports reliable decisions, and enables cumulative scientific understanding (Bazgir et al., 2025).
+To ground this discussion, consider spam detection as a binary classification task. A predicate-based rule set might include three predicate statements:
+1. $\mathrm{HasToken}(\texttt{prize},x)\wedge\mathrm{HasFile}(x)\rightarrow\mathrm{Spam}(x)$;
+2. $\mathrm{InBlacklist}(x)\wedge\mathrm{HastLink}(x)\rightarrow\mathrm{Spam}(x)$;
+3. $\mathrm{TitleAllCaps}(x)\wedge\mathrm{HasToken}(\texttt{free},x)\rightarrow\mathrm{Spam}(x)$.
+A deterministic inference framework is to predict spam if any of the conditions is satisfied (Fürnkranz and Kliegr, 2015).
+Such classical rule learning methods depend on a predefined predicate space with limited expressiveness for unstructured inputs and open-ended semantics (Cerna and Cropper, 2024; Hocquette et al., 2024).
+In contrast, probabilistic approaches treat each rule’s firing as an indicator ($r_{j}\in\{0,1\}$) and learn global weights – e.g., via logistic regression – to combine them into a unified prediction (Ruczinski et al., 2003; Friedman and Popescu, 2008).
+Crucially, this probabilistic formulation allows for greater flexability, accommondating rules defined as either symbolic predicates or neural indicators.
+Figure 1: The benefit of explicit aggregation.
+While direct LLM prompting struggles to resolve conflicts among rules with varying semantic levels (Top), RLIE’s structured approach—combining rule-wise judging with a weighted logistic combiner—effectively integrates evidence to yield a reliable prediction (Bottom).
+The advent of large language models (LLMs) opens a new opportunity for learning expressive rules in natural languages: given a few examples and task context, LLMs can generate rules directly in natural language (Ellis, 2023; Singh et al., 2022).
+Such rules can articulate fine-grained conditions, relations, and exceptions that are cumbersome to enumerate with a fixed predicate space.
+However, existing LLM-based methods typically either iteratively optimize a single rule (Qiu et al., 2023) or generate multiple rules without considering their global interdependence.
+This often leads to misalignment between training and inference (Yang et al., 2023; Zhou et al., 2024; Yang et al., 2024b), leaving the challenge of learning a compact and coordinated rule set unresolved.
+More broadly, the probabilistic composition of natural-language rules remains under-studied: it is unclear whether LLMs can robustly apply rule weights for reliable aggregation.
+Motivated by our empirical findings on the limitations of LLM probabilistic reasoning (see Section 5.2), we propose a framework that combines the expressive power of LLMs with the calibrated composition of classical probabilistic models.
+See Figure 1 for an illustrative example.
+To realize this synergy, we propose RLIE, a unified framework integrating Rule generation, Logistic regression, Iterative refinement, and Evaluation.
+RLIE first prompts an LLM to generate a pool of candidate rules.
+To enable learning, it uses the LLM to produce instance-level judgments for each rule.
+A regularized logistic regression model is then fitted to learn probabilistic weights, acting as explicit importance scores for rule selection.
+RLIE iteratively improves the rule set by mining hard examples to guide the LLM in revising or extending rules.
+Finally, the Evaluation stage verifies the system by comparing the learned logistic combiner against prompt-based inference.
+Our contributions can be summarized as follow:
+1.
+We advocate a principled division of labor: LLMs handle semantic rule induction and local assessment, while statistical models provide globally calibrated composition, balancing semantic flexibility with inference robustness.
+2. We propose a hard-example-driven refinement mechanism that improves the rule set through reflective revision and targeted extension, promoting rule quality and diversity.
+3.
+We evaluate inference paradigms, finding that although LLMs excel at proposing interpretable rules, they struggle to effectively utilize rule weights for precise reasoning.
+This highlights a key deficiency in LLM reasoning and underscores the necessity of an external probabilistic combiner.
 2 Related Work
@@
 Rule Learning.
-Canonical approaches to rule learning include Inductive Logic Programming (ILP)(Cropper et al., 2022), which is centered on search and constraint optimization; association and pattern mining methods that rely on heuristics and pruning; and differentiable neuro-symbolic methods(Qiao et al., 2021; Glanois et al., 2022; Yang et al., 2024a) that couple symbolic rules with continuous representations.
-A common objective across these paradigms is to produce a set of human-readable and reusable discriminative rules from a structured candidate space.
-However, they are fundamentally dependent on a predefined predicate space.
+Classical approaches encompass Inductive Logic Programming (ILP) (Cropper et al., 2022), association rule mining, and differentiable neuro-symbolic methods (Qiao et al., 2021; Glanois et al., 2022; Yang et al., 2024a).
+Despite their methodological differences, these paradigms share a common objective: inducing human-readable, reusable discriminative rules within a structured hypothesis space.
+A fundamental limitation remains their reliance on a predefined predicate vocabulary, which restricts expressiveness and hinders applicability to unstructured inputs with open-ended semantics.
 Rule Organization and Reasoning.
-In the traditional paradigm, the interaction among rules is determined by their organization and the semantics of the inference process.
-One approach is the use of ordered rule lists or decision lists(Yang et al., 2017; Xu et al., 2024), which employ sequential matching to explicitly encode relationships of priority, exception, and default cases.
-This is well-suited for scenarios with clear sequential or override logic.
-Another approach involves unordered rule sets(Qiao et al., 2021; Yang & van Leeuwen, 2022), which support the parallel integration of multiple pieces of evidence, facilitating global trade-offs and model compression.
-Within the unordered setting, deterministic aggregation methods (e.g., logical OR/AND, thresholding, majority voting) are simple and intuitive but offer limited flexibility when faced with local conflicts, coverage gaps, or inter-rule dependencies.
-Consequently, many works have introduced probabilistic aggregation to model uncertainty and improve calibration.
-A common and interpretable implementation is the linear log-odds model, which treats the binary satisfaction of each rule $r_{j}(x)\in\{0,1\}$ on a sample $x$ as a feature and learns a weighted linear combination(Ruczinski et al., 2003):
-$\mathrm{Pr}(y=1\mid x)=\sigma(\beta_{0}+\sum_{j}\beta_{j}r_{j}(x)).$
-Our work builds upon this foundation.
-Traditional methods typically assume rules are expressed in a fixed, executable symbolic form, allowing for direct evaluation of their satisfaction.
-In contrast, our approach focuses on rules expressed in natural language, where satisfaction must be determined by an LLM.
-We adopt a two-level design: at the “local judgment” level, we leverage the expressive power and adaptability of LLMs to interpret individual rules; at the “global aggregation” level, we retain the robust and calibratable probabilistic linear combiner from the classical paradigm.
-This hybrid design captures the flexible semantics of natural language rules while ensuring the overall inference process remains transparent and robust.
+Inference semantics dictate how multiple rules interact to form a decision.
+One stream of research utilizes ordered rule lists (e.g., decision lists) to encode priority, exceptions, and defaults via sequential matching (Yang et al., 2017; Xu et al., 2024).
+Conversely, unordered rule sets integrate evidence in parallel, enabling global trade-offs and model compression (Qiao et al., 2021; Yang and van Leeuwen, 2022).
+Within unordered settings, deterministic aggregators (e.g., OR/AND, thresholding, voting) offer simplicity but often prove brittle under conflicts, coverage gaps, and inter-rule dependencies.
+Probabilistic aggregation mitigates these issues by modeling uncertainty and improving calibration.
+A prominent interpretable instantiation is the linear log-odds model, which treats rule satisfaction $r_{j}(x)\in\{0,1\}$ as a feature and learns a weighted combination (Ruczinski et al., 2003) $\Pr(y=1\mid x)=\sigma\!\left(\beta_{0}+\sum_{j}\beta_{j}r_{j}(x)\right).$
+Our work builds upon this foundation but departs from the standard assumption that rules exist in a fixed, executable symbolic form with directly computable satisfactions.
+Instead, we focus on natural-language rules, whose instance-level satisfactions must be judged (and may involve abstention) by LLMs.
+This distinction motivates a principled division of labor: employing LLMs for local rule interpretation and instance-level judgments, while retaining a logistic combiner for global selection and calibrated aggregation.
+The resulting hybrid inference preserves the semantic expressiveness of natural-language rules while ensuring the stability and interpretability of the overall decision process.
 2.2 LLM-based Rule Learning and Its Applications
 Rule Learning.
-Large Language Models (LLMs) can directly propose hypotheses in natural language.
-A representative approach involves a “single-hypothesis refinement” loop, where the model maintains only the current best hypothesis and iteratively improves its quality and interpretability through feedback from validation sets or external executors(Qiu et al., 2023).
-In contrast, other methods adopt a “multi-hypothesis set” approach.
-For example, HypoGeniC(Zhou et al., 2024) generates multiple candidate hypotheses from data samples, maintains a top-$k$ set, and uses training signals to filter and update this set, ultimately forming an interpretable hypothesis bank.
-These two strategies exemplify the trade-off between deep optimization of a single hypothesis and the parallel retention of multiple diverse hypotheses.
+Current LLM-based induction falls into two paradigms: single-hypothesis refinement (Qiu et al., 2023), which iteratively optimizes one rule via feedback, and multi-hypothesis methods like HypoGeniC (Zhou et al., 2024) that maintain and update a candidate pool.
+However, these methods typically treat hypotheses in isolation, reflecting a trade-off between the intensive refinement of a single rule and the retention of a diverse rule set.
 Rule Organization and Reasoning.
-Rules expressed in natural language are typically reasoned over by the LLM that generated them.
-Existing work often passes one or more rules as context to an LLM, demonstrating that such information can enhance its reasoning capabilities(Zhang et al., 2024; Zhou et al., 2024).
-However, as the rules become more numerous or complex, the stability and instruction-following fidelity of LLMs can become unreliable.
-The black-box nature of LLMs makes it challenging to analyze the organization and interaction of these rules.
-Our work is distinct from existing LLM-based rule learning approaches in that we are the first to explicitly combine LLMs with probabilistic methods to learn a set of weighted rules.
-This integration enables more principled sample selection during the iterative refinement stage and leads to more flexible and robust inference.
-To better analyze the synergy between this new rule paradigm and LLMs, we systematically compare direct inference using the probabilistic combiner with a hierarchy of strategies for injecting the rules, their weights, and the combiner’s predictions into an LLM.
-This allows us to provide empirical analysis and practical guidance on the trade-offs in accuracy and robustness.
+Existing work predominantly uses the LLM as the inference engine via rule-augmented prompting (Zhang et al., 2024; Zhou et al., 2024).
+While effective in specific settings, implicit prompt-based aggregation becomes unstable as rule sets expand or conflict.
+Moreover, this black-box approach hinders error attribution, calibration, and the analysis of rule interactions, motivating our move toward explicit probabilistic composition.
+In contrast, our work investigates explicit probabilistic composition for natural-language rules.
+We systematically compare direct probabilistic combination against prompt-based inference under progressively richer injected information (rules, weights, and prediction references), demonstrating the necessity of an external, calibrated aggregator.
 3 Method
-We consider a binary classification task on a labeled dataset $\mathcal{S}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}\in\mathcal{X}$ is a natural language text and $y_{i}\in\{0,1\}$ is its corresponding label.
-The dataset $\mathcal{S}$ is partitioned into a training set $\mathcal{S}_{\mathrm{tr}}$, a validation set $\mathcal{S}_{\mathrm{val}}$ and a test set $\mathcal{S}_{\mathrm{te}}$.
-To perform rule based reasoning, our objective is to learn a set of natural language rules $\mathcal{H}^{\star}=\{h_{1},\ldots,h_{m}\}$ and a set of weights $\theta$ that aggregate the rules to model the discriminative relationship between $x$ and $y$.
-Here, $m\leq H$ where $H$ is a predefined capacity limit, and $h_{i}$ is a rule in natural language.
-Our method contains four stages, and the whole procedure is illustrated in Figure 1:
-Figure 1: The pipeline of the RLIE framework.
-(1) Rule Generation: An LLM proposes an initial set of rules from a small sample of training data.
-(2) Logistic Regression: A regularized logistic regression model is trained on the full training set to learn probabilistic weights for the current rules.
-(3) Iterative Refinement: Prediction errors from the regression model are used to mine hard examples, which are then used to prompt the LLM to refine or generate new rules.
-This cycle continues until performance on a validation set converges.
-(4) Evaluation: The learnt rules as well as weights are evaluated with different inference strategies.
-Detailed explanations of different strategies are in Section 3.4.
-  1. 1.
-    Rule Generation: We initialize an empty rule set $\mathcal{H}:=\emptyset$.
-An LLM is prompted with a randomly selected subset of $\mathcal{S}_{\mathrm{tr}}$ to generate candidate rules, which are then filtered based on coverage to form the initial rule set $\mathcal{H}^{(1)}$.
-  2. 2.
-    Logistic Regression: To obtain the optimal weights to combinatorially use the generated rules, logistic regression is performed to learn the rule weights $\theta=(\bm{\beta},b)$ on $\mathcal{S}_{\mathrm{tr}}$ with hyperparameters tuned using $\mathcal{S}_{\mathrm{val}}$.
-  3. 3.
-    Iterative Refinement: Hard examples are selected from $\mathcal{S}_{\mathrm{tr}}$ based on prediction errors.
-These samples, along with the current rule set $\mathcal{H}^{(t)}$, are fed back to the LLM to generate a set of new rules to revise existing rules.
-The union of newly generated rules and old rules can be directly used as $\mathcal{H}^{(t+1)}$, if it doesn’t exceed the capacity limit $H$, otherwise we need to discard some rules based on their classification accuracy.
-The rule weights $\theta^{(t+1)}$ are also updated.
-This process is repeatedly executed for several times to iteratively refine the rules, while the whole process is monitored on $\mathcal{S}_{\mathrm{val}}$ for early stopping.
-  4. 4.
-    Evaluation: After getting generated rules, we conduct a comprehensive evaluation on $\mathcal{S}_{\mathrm{te}}$ using two classes of inference strategies: direct prediction using the linear combiner, and injecting the learnt rules and their weights into an LLM for augmented reasoning, to evaluate the quality of rules with different inference strategies.
+We consider a binary classification task on a labeled dataset $\mathcal{S}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}\in\mathcal{X}$ is a natural-language text and $y_{i}\in\{0,1\}$ is its label.
+We split $\mathcal{S}$ into training, validation, and test sets, denoted by $\mathcal{S}_{\mathrm{train}}$, $\mathcal{S}_{\mathrm{val}}$, and $\mathcal{S}_{\mathrm{test}}$, respectively.
+Our goal is to learn a set of natural-language rules $\mathcal{H}^{\star}=\{h_{1},\ldots,h_{m}\}$ together with aggregation weights $\theta$ that linearly combine these rules to model the discriminative relation between $x$ and $y$.
+We constrain the rule set size by $m\leq H$, where $H$ is a predefined capacity limit.
+Our framework consists of four stages (Figure 2):
+Figure 2: Overview of the RLIE pipeline.
+(1) Rule Generation: An LLM proposes candidate rules from a small sample of $\mathcal{S}_{\mathrm{train}}$, which are then filtered by coverage to form an initial rule set.
+(2) Weight Learning: A regularized logistic regression combiner is trained on $\mathcal{S}_{\mathrm{train}}$ to learn rule weights.
+(3) Iterative Refinement: Hard samples are mined based on the combiner’s errors to prompt the LLM to revise or add rules; the rule set is pruned to satisfy a capacity limit and training proceeds with early stopping on $\mathcal{S}_{\mathrm{val}}$.
+(4) Evaluation: The learned rules and weights are assessed under different inference protocols.
+1.
+Rule Generation: Starting from $\mathcal{H}:=\emptyset$, we prompt an LLM with a randomly sampled subset of $\mathcal{S}_{\mathrm{train}}$ to propose candidate rules, and filter them by coverage to obtain the initial rule set $\mathcal{H}^{(1)}$.
+2.
+Logistic Regression: Given $\mathcal{H}^{(t)}$, we train a logistic regression combiner on the whole training set $\mathcal{S}_{\mathrm{train}}$ to learn rule weights $\theta^{(t)}$, with hyperparameters selected on $\mathcal{S}_{\mathrm{val}}$.
+3.
+Iterative Refinement: We select hard examples from $\mathcal{S}_{\mathrm{train}}$ with the highest prediction errors from the combiner, and feed them back to the LLM together with the current rule set $\mathcal{H}^{(t)}$ to generate new rules $\Delta\mathcal{H}^{(t)}$ to augment the rule set.
+We then merge the updated rules and enforce the capacity constraint: if $|\mathcal{H}^{(t)}\cup\Delta\mathcal{H}^{(t)}|>H$, we prune rules based on validation performance.
+The combiner is retrained to update $\theta^{(t+1)}$.
+We repeat this process with early stopping monitored on $\mathcal{S}_{\mathrm{val}}$.
+4.
+Evaluation: We strictly evaluate the finalized rule set and weights on $\mathcal{S}_{\mathrm{test}}$.
+This stage quantifies the system’s reliability and compares the calibrated probabilistic inference (combiner) against uncalibrated LLM reasoning (prompting), verifying the effectiveness of weighted rule set.
+Our prompts are provided in Appendix E.
 3.1 Rule Generation
 Initialization and Candidate Generation.
-In the first iteration, we initialize the rule set as empty, $\mathcal{H}^{(0)}:=\emptyset$.
-We randomly sample $k$ examples from $\mathcal{S}_{\mathrm{tr}}$ and prompt an LLM to generate $h$ candidate rules (where $h\leq H$), denoted as $\mathcal{H}_{\mathrm{cand}}^{(1)}=\{{h}^{(1)}_{1},\ldots,{h}_{h}^{(1)}\}$.
-Individual Rule Application. For any samples $x_{i}$ and candidate rule ${h}_{j}^{(1)}$, we prompt an LLM to obtain a local, ternary judgment
-$z_{i,j}^{(1)}=\mathrm{LLM}\left(x_{i},{h}_{j}^{(i)}\right)\in\{-1,0,+1\},$
-as $z_{i,j}$ means the evaluation results of applying the $j$-th rule on the $i$-th sample where +1+1 indicates a positive prediction, 1-1 a negative prediction and 0 signifies the rule is not applicable (“abstain”).
-Allowing abstention is crucial for explicitly modeling rule coverage, reducing misclassifications, and enabling sparse, robust combinations in the subsequent stage.
-Coverage-based Filtering and Rule Set Update. For each candidate rule ${h}_{j}^{(1)}$, we compute its coverage on the training set as
-$\mathrm{Cov}_{\mathrm{tr}}\left({h}_{j}^{(1)}\right)=\frac{1}{N_{\mathrm{tr}}}\sum_{i=1}^{N_{\mathrm{tr}}}\mathbb{I}\left[z_{i,j}^{(1)}\neq 0\right],$
-and discard rules with coverage lower than a predefined threshold $\gamma$.
-After $\mathrm{Cov}_{\mathrm{tr}}$ based filtering, the valid rules generated here is denoted as $\mathcal{H}_{\mathrm{new}}^{(1)}$, and the rule set is updated as $\mathcal{H}^{(1)}=\mathcal{H}^{(0)}\cup\mathcal{H}_{\mathrm{new}}^{(1)}$.
+We start with an empty rule set $\mathcal{H}^{(0)}:=\emptyset$.
+In the first iteration, we uniformly sample $k$ examples from $\mathcal{S}_{\mathrm{train}}$ and prompt an LLM to propose $h$ candidate natural-language rules ($h\leq H$), denoted by $\mathcal{H}_{\mathrm{cand}}^{(1)}=\{h^{(1)}_{1},\ldots,h^{(1)}_{h}\}$.
+Individual Rule Application.
+For each candidate rule $h_{j}^{(1)}$ and sample $x_{i}$ from the whole training set, we query an LLM-based judge to produce a ternary judgment $z_{i,j}^{(1)}=\mathrm{LLM}\!\left(x_{i},\,h_{j}^{(1)}\right)\in\{-1,0,+1\},$ where +1+1 and 1-1 indicate predictions for the positive and negative class, respectively, and 0 denotes abstention when the LLM deems the rule inapplicable to sample $x_{i}$.
+This explicitly models rule coverage and mitigates spurious forced predictions, facilitating sparse and robust aggregation in the subsequent stage.
+Coverage-based Filtering and Rule Set Update.
+We measure the coverage of each candidate rule $h_{j}^{(1)}$ on the training set as: $\mathrm{Cov}_{\mathrm{train}}\!\left(h_{j}^{(1)}\right)=\frac{1}{N_{\mathrm{train}}}\sum_{i=1}^{N_{\mathrm{train}}}\mathbb{I}\!\left[z_{i,j}^{(1)}\neq 0\right].$ We discard rules from $\mathcal{H}_{\mathrm{cand}}^{(1)}$ with $\mathrm{Cov}_{\mathrm{train}}\!\left(h_{j}^{(1)}\right)<\gamma$ to ensure that the rule set is general and applicable to a significant portion of samples.
+We denote the remaining rules by $\mathcal{H}_{\mathrm{new}}^{(1)}$.
+The rule set is then updated as $\mathcal{H}^{(1)}=\mathcal{H}^{(0)}\cup\mathcal{H}_{\mathrm{new}}^{(1)}$.
 3.2 Logistic Regression
+Unlike neural networks which are sometimes overconfident(Guo et al., 2017), our logistic combiner minimizes log-loss, inherently promoting calibration.
 Feature Construction.
-For any iteration $t$, given a rule set $\mathcal{H}^{(t)}=\{h_{1}^{(t)},\ldots,h^{(t)}_{m^{(t)}}\}$ where $m^{(t)}\leq H$, we define a mapping $\Phi^{(t)}$ from an input sample to a vector of rule judgments
-$\Phi^{(t)}:\ \mathcal{X}\to\{-1,0,+1\}^{m^{(t)}},\qquad\Phi^{(t)}(x_{i})=\mathbf{z}_{i}^{(t)}=\begin{bmatrix}z_{i,1}^{(t)},\ldots,z_{i,m^{(t)}}^{(t)}\end{bmatrix}^{\top}$.
-Weight Learning. The probability of a sample $x_{i}$ being classified as positive is modeled by a logistic regression function, which is defined as below
-$p^{(t)}(x_{i};\theta^{(t)})=\sigma\!\big((\Phi^{(t)}(x_{i}))^{\top}\bm{\beta}^{(t)}+b^{(t)}\big),\qquad\sigma(u)=(1+e^{-u})^{-1}$
-where $\sigma$ is the sigmoid function and $\theta^{(t)}=(\bm{\beta}^{(t)},b^{(t)})$ are the model parameters.
-We learn these parameters by minimizing the cross-entropy loss on $\mathcal{S}_{\mathrm{tr}}$ with Elastic Net (Zou & Hastie, 2005) regularization:
-$\bm{\beta}^{(t)},b^{(t)}=\underset{{\bm{\beta}^{(t)},b^{(t)}}}{\arg\min}\ \frac{1}{|\mathcal{S}_{\mathrm{tr}}|}\!\sum_{x_{i}\in\mathcal{S}_{\mathrm{tr}}}\mathcal{L}(y_{i},p^{(t)}(x_{i};\theta^{(t)}))+\lambda\!\left(\alpha\|\bm{\beta}^{(t)}\|_{1}+\tfrac{1-\alpha}{2}\|\bm{\beta}^{(t)}\|_{2}^{2}\right),$
-where $\mathcal{L}$ is the binary cross-entropy loss.
-The $L_{1}$ penalty encourages sparsity (rule selection), while the $L_{2}$ penalty enhances robustness.
-The regularization hyperparameters $(\lambda,\alpha)$ are selected via stratified K-fold cross-validation on $\mathcal{S}_{\mathrm{val}}$.
-The final parameters $\hat{\theta}^{(t)}=(\hat{\bm{\beta}}^{(t)},\hat{b}^{(t)})$ are obtained by refitting the model on the entire $\mathcal{S}_{\mathrm{tr}}$ with the selected hyperparameters.
-Label Prediction. After getting rule weights, the prediction for a sample $x_{i}$ can be calculated as
-$\hat{y}_{i}^{(t)}=\mathbb{I}\big[p^{(t)}(x_{i};\hat{\theta}^{(t)})\geq\tau\big]\in\{0,1\},\qquad\text{with }\tau=0.5.$
+At iteration $t$, given a rule set $\mathcal{H}^{(t)}=\{h_{1}^{(t)},\ldots,h^{(t)}_{m^{(t)}}\}$ where $m^{(t)}\leq H$, we define a mapping $\Phi^{(t)}$ from an input sample to a vector of rule application results $\Phi^{(t)}:\mathcal{X}\to\{-1,0,1\}^{m^{(t)}},\quad\Phi^{(t)}(x_{i})=\mathbf{z}_{i}^{(t)}=[z_{i,j}^{(t)}]_{j=1}^{m^{(t)}}.$
+Weight Learning.
+We model the positive-class probability using a regularized logistic regression combiner $p^{(t)}(x_{i};\theta^{(t)})=\sigma\!\big((\Phi^{(t)}(x_{i}))^{\top}\boldsymbol{\beta}^{(t)}+b^{(t)}\big)$ where $\sigma$ is the sigmoid function and $\theta^{(t)}=(\boldsymbol{\beta}^{(t)},b^{(t)})$.
+We learn these parameters by minimizing the cross-entropy (CE) loss on $\mathcal{S}_{\mathrm{train}}$ with Elastic Net (Zou and Hastie, 2005) regularization:
+$\begin{split}\boldsymbol{\beta}^{(t)},\,b^{(t)}=\mathop{\arg\min}_{\boldsymbol{\beta}^{(t)},\,b^{(t)}}\Bigg[&\frac{1}{|\mathcal{S}_{\mathrm{train}}|}\sum_{x_{i}\in\mathcal{S}_{\mathrm{train}}}\mathcal{L}\!\left(y_{i},\,p^{(t)}(x_{i};\theta^{(t)})\right)\\ &+\lambda\left(\alpha\|\boldsymbol{\beta}^{(t)}\|_{1}+\frac{1-\alpha}{2}\|\boldsymbol{\beta}^{(t)}\|_{2}^{2}\right)\Bigg]\end{split}$
+where $\mathcal{L}$ denotes binary cross-entropy.
+The $L_{1}$ term promotes sparsity (rule selection), while the $L_{2}$ term improves robustness.
+We select $(\lambda,\alpha)$ on $\mathcal{S}_{\mathrm{val}}$ via stratified $K$-fold cross-validation, and refit on the full $\mathcal{S}_{\mathrm{train}}$ to obtain $\hat{\theta}^{(t)}$.
+Label Prediction. Given $\hat{\theta}^{(t)}$, we predict sample $x_{i}$ with $\hat{y}_{i}^{(t)}=\mathbb{I}\big[p^{(t)}(x_{i};\hat{\theta}^{(t)})\geq 0.5\big]\in\{0,1\}.$
 3.3 Iterative Refinement
-In the first iteration, $\mathcal{H}^{(1)}$ and $\hat{\theta}^{(1)}$ are obtained on randomly selected samples.
-In the subsequent iterations, given $\mathcal{H}^{(t)}$ and $\hat{\theta}^{(t)}$ with $t\geq 1$, to generate $\mathcal{H}^{(t+1)}$ and $\hat{\theta}^{(t+1)}$, instead of relying on randomly selected samples, we can enhance the rule generation by targeting hard examples.
+Although the initial rule set $\mathcal{H}^{(1)}$ is derived from a random subset of $\mathcal{S}_{\mathrm{train}}$, we subsequently refine it by targeting hard examples under the current model $(\mathcal{H}^{(t)},\hat{\theta}^{(t)})$ for $t\geq 1$.
 1.
 Hard Example Selection.
-For each example $x_{i}\in\mathcal{S}_{\mathrm{tr}}$, we compute its predicted probability $\hat{p}_{i}^{(t)}=p^{(t)}(x_{i};\hat{\theta}^{(t)})$ and define its prediction error as $d_{i}=\|\hat{p}_{i}^{(t)}-y_{i}\|$.
-We identify the top-$k$ samples with the highest error $I_{\mathrm{hard}}^{(t)}=\mathrm{TopK}\left(\{(i,d_{i})\}^{\|\mathcal{S}_{\mathrm{tr}}\|}_{i=1},k\right)$.
+For each $(x_{i},y_{i})\in\mathcal{S}_{\mathrm{train}}$, we compute the predicted probability $\hat{p}_{i}^{(t)}=p^{(t)}(x_{i};\hat{\theta}^{(t)})$ and define the prediction error $d_{i}=|\hat{p}_{i}^{(t)}-y_{i}|$.
+We then select the indices of the top-$k$ hardest examples: $\mathcal{I}_{\mathrm{hard}}^{(t)}=\mathrm{TopK}\!\left(\{d_{i}\}_{i=1}^{|\mathcal{S}_{\mathrm{train}}|},\,k\right).$
 2.
 New Rule Generation.
-The selected hard examples $I^{(t)}_{\mathrm{hard}}$ are presented to the LLM along with the current rule set $\mathcal{H}^{(t)}$.
-The LLM is prompted to reflect on the errors and either revise existing rules or generate new ones that cover novel perspectives.
-The $h$ newly generated rules are filtered for coverage, resulting in the set $\mathcal{H}_{\mathrm{new}}^{(t+1)}$.
+We present the hard examples $\mathcal{I}_{\mathrm{hard}}^{(t)}$ along with the current rule set $\mathcal{H}^{(t)}$ to the LLM, prompting it to revise existing rules or propose $h$ new rules (see Figure 6).
+After applying the same coverage-based filtering as in Section 3.1, we obtain $\mathcal{H}_{\mathrm{new}}^{(t+1)}$.
 3.
 Rule Set Update.
-The new rules are merged with the existing set: $\mathcal{H}_{\mathrm{tmp}}^{(t+1)}=\mathcal{H}^{(t)}\cup\mathcal{H}_{\mathrm{new}}^{(t+1)}$.
-If the size of this temporary set does not exceed the capacity $H$, it can be directly used as $\mathcal{H}^{(t+1)}$, otherwise we have to prune the rules by ranking them based on their individual accuracy on the validation set, retaining only the top $H$ rules to form $\mathcal{H}^{(t+1)}$.
+We merge new rules with the current set: $\mathcal{H}_{\mathrm{tmp}}^{(t+1)}=\mathcal{H}^{(t)}\cup\mathcal{H}_{\mathrm{new}}^{(t+1)}$.
+If $|\mathcal{H}_{\mathrm{tmp}}^{(t+1)}|\leq H$, we set $\mathcal{H}^{(t+1)}=\mathcal{H}_{\mathrm{tmp}}^{(t+1)}$.
+Otherwise, we prune rules by ranking them according to their F1 scores on $\mathcal{S}_{\mathrm{val}}$ calculated on non-abstaining samples and retaining the top $H$ rules.
 4.
 Parameter Update and Termination.
-With the updated rule set $\mathcal{H}^{(t+1)}$, we calculate $\hat{\theta}^{(t+1)}$ in the same way as discussed in Section 3.2, after which we can monitor the overall performance on $\mathcal{S}_{\mathrm{val}}$.
-Rules will be refined iteratively, with each iteration trying to improve the performance on hard examples of previous version of rule sets.
-The process terminates if the overall performance on $\mathcal{S}_{\mathrm{val}}$ fails to improve by a margin $\delta$ for $p$ consecutive iterations or if the maximum number of iterations $R_{\mathrm{max}}$ is reached.
-The final model $(\mathcal{H}^{\star},\hat{\theta}^{\star})$ corresponds to the checkpoint with the best validation performance.
+Given $\mathcal{H}^{(t+1)}$, we refit the logistic regression combiner as described in Section 3.2 to obtain $\hat{\theta}^{(t+1)}$.
+The refinement terminates if validation performance fails to improve by at least $\delta$ for $p$ consecutive iterations, or if the maximum iteration count $R_{\mathrm{max}}$ is reached.
+We return $(\mathcal{H}^{\star},\hat{\theta}^{\star})$ from the checkpoint with the best validation performance.
 3.4 Evaluation
-We systematically compare four inference strategies on the test set $\mathcal{S}_{\mathrm{te}}$, categorized into two main approaches: direct inference using the linear combiner and LLM-augmented inference with varying levels of information.
-(E1) Linear-only.
-The final prediction is made directly by the logistic regression model, with $\hat{y}_{i}^{\mathrm{final}}=\mathbb{I}\big[p(x_{i};\hat{\theta}^{\star})\geq 0.5]$, where $p(\cdot;\hat{\theta}^{\star})$ is the learnt probabilistic combiner.
-This serves as a classic baseline for weighted rule sets.
-(E2) LLM + Rules.
-The LLM is provided with only the texts of the rules in $\mathcal{H}^{\star}$ and the input $x_{i}$, and is tasked with making a prediction where $\hat{y}_{i}^{\mathrm{final}}=\mathrm{LLM}(x_{i},\mathcal{H}^{\star})$.
-This evaluates the LLM’s intrinsic ability to aggregate rule-based evidence, mirroring common practices in existing work.
-(E3) LLM + Rules + Weights.
-In addition to the rule texts, the LLM is also provided with the learnt weights $\hat{\theta}^{\star}$ and the final prediction becomes $\hat{y}_{i}^{\mathrm{final}}=\mathrm{LLM}(x_{i},\mathcal{H}^{\star},\hat{\theta}^{\star})$.
-This tests whether the LLM can leverage explicit probabilistic signals to guide its reasoning, reflecting a “probabilistic internalization” of rule importance.
-(E4) LLM + Rules + Weights + Linear Prediction.
-The LLM receives all information from generated rules and corresponding weights, plus the prediction $\hat{y}_{i}$ from the linear-only model as a reference, so the prediction becomes $\hat{y}_{i}^{\mathrm{final}}=\mathrm{LLM}(x_{i},\mathcal{H}^{\star},\hat{\theta}^{\star},\hat{y}_{i})$.
-This strategy assesses whether a potentially imperfect but calibrated rule-based system can provide a valuable external signal to enhance the LLM’s reasoning in complex scenarios where linear assumption may not fully hold.
+Upon obtaining the weighted rule sets, we employ various inference strategies to verify their quality and identify the optimal deployment protocol.
+These strategies fall into two paradigms: direct inference using the learned logistic combiner, and LLM-based inference where the prompt is progressively augmented with information from the learned system.
+Let $\mathrm{LLM}(\cdot)$ denote an LLM that outputs the final label.
+We use $\mathbf{R}$ to denote the rule texts in $\mathcal{H}^{\star}$, $\mathbf{W}$ the learned weights $\hat{\theta}^{\star}$, and $\mathbf{P}$ the combiner’s predicted label $\hat{y}_{i}^{\mathrm{comb}}=\mathbb{I}\!\left[p(x_{i};\hat{\theta}^{\star})\geq 0.5\right]$.
+(E1) Combiner. We set $\hat{y}_{i}^{\mathrm{final}}=\hat{y}_{i}^{\mathrm{comb}}$, directly utilizing the combiner’s output as the final prediction without further LLM prompting at inference time.
+(E2) LLM(R).
+The LLM is provided with the input $x_{i}$ and rule texts $\mathbf{R}$: $\hat{y}_{i}^{\mathrm{final}}=\mathrm{LLM}(x_{i},\mathbf{R})$.
+This tests the LLM’s intrinsic ability to reason over a set of natural-language rules absent of explicit weight signals.
+(E3) LLM(RW).
+The LLM is provided with $x_{i}$, $\mathbf{R}$, and the learned weights $\mathbf{W}$: $\hat{y}_{i}^{\mathrm{final}}=\mathrm{LLM}(x_{i},\mathbf{R},\mathbf{W})$.
+This tests whether explicit probabilistic importance cues facilitate the LLM in prioritizing and aggregating rule evidence.
+(E4) LLM(RWP).
+The LLM is provided with $x_{i}$, $\mathbf{R}$, $\mathbf{W}$, and the combiner prediction $\mathbf{P}$: $\hat{y}_{i}^{\mathrm{final}}=\mathrm{LLM}(x_{i},\mathbf{R},\mathbf{W},\mathbf{P})$.
+This tests whether the combiner’s prediction can serve as a reference signal to stabilize LLM-based aggregation in challenging cases.
 4 Experiment Setup
 4.1 Dataset
-To evaluate our method, we select six real-world tasks from the HypoBench(Liu et al., 2025) benchmark, including Deception Detection (Review), Persuasive Argument Prediction (Dreddit), News Headline Engagement (Headlines), Paper Citations (Citations), AI-generated Content Detection (LLM Detect) and Retweets (Retweets).
-Each of the task can be viewed as a binary classification task, with text information as inputs and binary classification results as output (e.g.
+We evaluate RLIE on six real-world tasks, including Reviews (deceptive review detection), Dreaddit (mental stress detection), Headlines (headline engagement prediction), Citations (paper citation impact prediction), LLM Detect (AI-generated text detection), and Retweets (pairwise retweet prediction).
+All tasks are formulated as binary classification over natural-language inputs (e.g.
 given a review, we need to evaluate whether it is deceptive or not).
-Details of each dataset are in Appendix A.3.
+Detailed dataset descriptions are provided in Appendis A.3.
+Table 1: Overall performance comparison (Accuracy / Macro-F1, $\times 10^{-2}$) on full datasets.
+RLIE (Ours) uses the Linear-Only inference strategy.
+The best results among generalizable methods are bolded.
+Method Backbone Reviews Dreaddit Headlines Citations LLM Detect Retweets
+Zero-shot DeepSeek-V3.2 53.9 / 42.3 64.7 / 60.0 61.2 / 59.6 62.5 / 50.0 82.4 / 82.0 63.5 / 62.6
+Few-shot (ICL) DeepSeek-V3.2 65.3 / 65.1 63.8 / 58.5 62.0 / 61.8 60.4 / 49.8 80.4 / 79.6 57.7 / 53.9
+Zero-shot Gen DeepSeek-V3.2 65.4 / 64.8 67.2 / 63.9 57.5 / 57.4 46.1 / 47.7 63.1 / 57.6 58.1 / 58.1
+IO Refinement DeepSeek-V3.2 65.9 / 65.4 78.5 / 78.1 62.0 / 61.1 54.2 / 51.0 83.6 / 83.4 57.1 / 56.1
+HypoGeniC DeepSeek-V3.2 69.1 / 69.3 80.5 / 80.5 59.9 / 60.1 46.9 / 49.3 85.2 / 85.1 61.9 / 61.8
+RLIE (Ours) Qwen3-Next-80B 68.3 / 67.8 81.1 / 81.1 61.1 / 60.9 56.3 / 55.0 87.6 / 87.6 61.9 / 61.8
+RLIE (Ours) Qwen3-235B 71.5 / 71.4 79.9 / 79.9 60.6 / 60.4 61.5 / 60.6 88.3 / 88.3 66.5 / 66.5
+RLIE (Ours) DeepSeek-V3.2 70.9 / 70.7 82.3 / 82.3 67.0 / 67.0 64.6 / 63.0 90.7 / 90.7 65.7 / 65.6
 4.2 Baseline
-We implemented and compared RLIE against several representative and competitive methods within a unified framework to comprehensively evaluate its performance and robustness.
-(1) Zero-shot Inference: The LLM directly predicts the output for a given sample based only on the task description, without access to any examples or rules.
-(2) Zero-shot Generation: The LLM generates a set of candidate rules based solely on the task description.
-The single best-performing rule, as determined on the validation set, is then used for inference on the test set.
-(3) IO Refinement (Qiu et al., 2023): This method follows a “propose-select-refine” loop.
-In each round, the LLM generates a batch of candidate hypotheses, the best one is selected based on its performance on a validation or error set, and it is then used as a seed for the next round of refinement.
-This process is repeated for a fixed number of iterations.
-(4) HypoGeniC (Zhou et al., 2024): This approach maintains and dynamically expands a library of hypotheses.
-For each sample, the top-$k$ hypotheses are selected to make a prediction.
-Mismatched samples are added to an error set, which triggers the generation of new hypotheses to improve coverage.
-The hypothesis library is continuously updated based on a reward signal.
+We compare RLIE with representative LLM-based inference and rule-learning baselines: (1) Zero-shot: direct prediction from task description without labeled demonstrations or rules.
+(2) Few-shot (ICL): standard in-context learning with a small set of labeled demonstrations sampled from the training split.
+(3) Zero-shot Gen: generate a pool of candidate rules from task description only, then select the single best rule on the validation set for test-time inference.
+(4) IO Refinement (Qiu et al., 2023): iteratively propose-select-refine a single rule using an error-driven refinement loop.
+(5) HypoGeniC (Zhou et al., 2024): maintain and expand a hypothesis library, and perform top-$k$ hypothesis selection for prediction with reward-driven updates.
 4.3 Experimental Details
-In our experiments, each dataset was partitioned into fixed-size training, validation, and test sets of 200, 200, and 300 samples, respectively.
-All methods used the same data splits and a fixed random seed to ensure reproducibility.
-We report Accuracy and Macro-F1 score.
-Each experiment was repeated at least three times, and we report the mean and standard deviation of the results.
-All experiments involving LLMs utilized gpt-4o-mini with the temperature set to $1\times 10^{-5}$ to ensure deterministic outputs.
-Prompt details can be found in Appendix C.
-As for hyperparameters, the capacity of the rule set was set to $H=10$.
+LLM backbones.
+To ensure a fair comparison for all methods that generate or refine rules, we use DeepSeek-V3.2 as the default backbone throughout the main experiments.
+To examine whether our rule-generation framework is model-agnostic, we additionally run RLIE with alternative backbones (Qwen3-Next-80B and Qwen3-235B) while keeping the rest of the pipeline unchanged.
+Protocol and metrics.
+We use the original train/validation/test splits provided by HypoBench for all methods, and report mean Accuracy (ACC) and Macro-F1 on the test set over three runs with different random seeds.
+RLIE hyperparameters.
+We set the rule capacity to $H=10$.
 In each iteration, we sampled $k=20$ hard examples and generated $h=5$ new rules.
-The minimum coverage threshold for rules was set to $\gamma=0.2$.
-The implementations of IO Refinement and HypoGeniC followed the default hyperparameter settings of the original papers.
+We filter candidate rules by a minimum training coverage threshold $\gamma=0.2$.
+During iterative refinement, we set the max iteration number $R_{\mathrm{max}}$ to be $20$.
+The sampling temperature is set to be $1\times 10^{-5}$ for stability and reproducibility.
+The implementations of IO Refinement and HypoGeniC follow the original papers.
 Further details are provided in Appendix A.
 5 Results
-This section presents empirical evaluation of our framework.
-First, we compare the overall performance of RLIE against competitive baselines.
-Second, we conduct an ablation study on different inference strategies to identify the most effective way to utilize the learned probabilistic rules.
-Table 1: Overall performance of different frameworks evaluated in Accuracy and F1 Score ($mean\pm std$, $\times 10^{-2}$).
-The best-performing results are shown in bold, while the second best results are highlighted in underline.
-The performance of RLIE is evaluated with the Linear Only strategy.
-Method Reviews Dreaddit Headlines Citations LLM Detect Retweets
-Accuracy
-Zero-shot Inference 55.11±0.18 62.00±0.88 63.22±2.52 56.25±0.00 56.11±0.38 65.56±2.34
-Zero-shot Generation 63.89±5.50 64.33±2.19 53.67±1.20 58.33±3.61 52.44±1.07 57.67±2.40
-IO Refinement 64.11±5.52 74.89±3.47 59.67±3.51 48.96±1.80 60.33±17.32 62.44±0.69
-HypoGeniC 58.00±6.35 65.67±0.00 60.78±3.27 47.92±7.22 54.56±4.43 65.56±2.12
-RLIE (Ours) 71.00±0.67 69.11±1.17 65.56±1.50 59.38±3.12 58.67±2.65 65.78±1.71
-F1 Score
-Zero-shot Inference 50.36±2.30 55.67±2.30 62.70±2.53 36.00±0.00 46.77±0.68 64.69±2.20
-Zero-shot Generation 63.41±6.09 59.17±2.84 53.10±0.66 42.56±7.04 39.68±1.58 57.64±2.42
-IO Refinement 62.03±8.45 73.98±3.59 59.21±3.23 45.66±2.19 54.00±22.42 62.30±0.66
-HypoGeniC 56.82±6.29 60.57±0.61 59.56±3.82 41.96±3.40 44.17±8.75 65.49±2.17
-RLIE (Ours) 70.81±0.59 65.94±1.25 65.38±1.44 49.08±2.84 52.10±4.51 65.45±1.41
+In this section, we analyse the overall performance of RLIE, conducting experiments on different inference strategies to find best way to apply the learnt rules.
+We also provide computational cost analysis and two case studies.
 5.1 Main Results
-The empirical results, summarized in Table 1, demonstrate the effectiveness and robustness of RLIE compared to several strong baselines across six real-world datasets.
-For this comparison, RLIE’s performance is evaluated using its direct inference strategy (Linear-only), where the learned logistic regression model acts as the classifier.
-While some baseline methods may attain the top performance on one or two individual datasets, RLIE consistently ranks within the top two in terms of both Accuracy and F1 Score.
-This demonstrates that the RLIE framework can be generalized to diverse data distributions and is effective at capturing underlying data patterns.
-As shown in the results table, some methods exhibit high variance and instability despite their strong performance on certain tasks (e.g., IO Refinement).
-In contrast, our method achieves high performance while maintaining stability, underscoring its robustness.
-A horizontal comparison of the different methods reveals several interesting insights.
-It is noteworthy that Zero-shot Inference outperforms many of the more complex baselines in several scenarios.
-This may be attributed to the extensive knowledge acquired by the LLM during pre-training, which is sufficient for solving certain tasks.
-However, this internal inference process is inherently unexplainable, and it is difficult to rule out the influence of potential data leakage from the pre-training corpus.
-This observation highlights both the challenge and the necessity of using rule-based learning to achieve interpretable reasoning.
-HypoGeniC learns a set of rules independently and fails to account for the interactions between them, resulting in suboptimal performance.
-By iteratively refining a single rule, IO Refinement achieves respectable results, which validates the value of providing annotated examples and employing an iterative optimization process.
-It is worth mentioning that in some cases, IO Refinement outperforms RLIE.
-This could be because the strategy of generating only a single rule forces it to be more generalizable, allowing it to better handle unseen scenarios.
-However, this single-rule approach also limits the method’s expressiveness, leading to poor performance in cases that require a more complex rule set.
-Our RLIE framework strikes a balance: by continuously learning from hard examples, it enhances the generalizability of the learned rule set, while the strategy of maintaining multiple rules significantly boosts its expressiveness.
-This allows our method to achieve excellent and robust performance with low variance across diverse datasets.
-Table 2: Performance of different inference strategies valuated in Accuracy and F1 Score ($mean\pm std$, $\times 10^{-2}$).
-The best-performing results are shown in bold, while the second best results are highlighted in underline.
-Inference Strategy Reviews Dreaddit Headlines Citations LLM Detect Retweets
-Accuracy
-Regression Only 71.00±0.67 69.11±1.17 65.56±1.50 59.38±3.12 58.67±2.65 65.78±1.71
-LLM Evaluation 65.22±2.67 64.56±0.51 61.33±3.06 60.42±1.80 55.22±2.12 61.00±3.06
-LLM + Linear Regression 64.67±2.33 67.56±2.50 61.89±3.02 60.42±1.80 54.11±2.22 62.22±3.56
-LLM + Linear + Label 64.67±2.60 69.00±1.33 65.33±1.86 59.38±3.12 56.67±3.51 60.22±3.56
-F1 Score
-Regression Only 70.81±0.59 65.94±1.25 65.38±1.44 49.08±2.84 52.10±4.51 65.45±1.41
-LLM Evaluation 64.52±3.19 59.57±0.94 60.88±3.18 48.62±1.19 46.82±5.32 60.33±3.13
-LLM + Linear Regression 63.87±3.05 63.73±3.72 61.71±2.93 47.11±3.38 44.76±4.62 59.68±4.45
-LLM + Linear + Label 63.80±2.87 65.79±1.51 65.16±1.78 49.08±2.84 48.70±6.05 57.38±4.84
+To enable a direct comparison with baselines, we evaluate RLIE using the combiner for final prediction throughout this section.
+The results are summarized in Table 1, showing that RLIE is strongly competitive across all tasks.
+We further evaluate RLIE with three different backbones and observe a consistent trend: stronger backbones typically yield a higher overall ceiling, while even with a smaller backbone RLIE remains highly competitive on multiple tasks, indicating robust adaptability across different model capacity regimes.
+Comparing against interpretable rule-based baselines reveals a clear progression in capability.
+Zero-shot Gen generates a set of candidate rules and selects the best single rule for inference, which is limited in expressiveness as it relies on a single hypothesis.
+IO Refinement iteratively improves a single rule, typically leading to stable gains by enhancing rule quality.
+HypoGeniC extends this to a set of multiple rules, which can provide additional benefits in some cases but does not consistently outperform single-rule methods, suggesting that multi-rule inference requires a more reliable global integration mechanism.
+Building on these observations, RLIE couples iterative rule expansion with weight learning and global aggregation: it progressively addresses uncovered hard cases while explicitly learning rule contributions and combining them at the global level, thereby leveraging the benefits of rule sets more reliably and achieving consistent improvements across most tasks.
+We also consider direct prompting (Zero-shot/Few-shot) in which LLMs directly output judgments without producing explicit rules, thus offering weaker interpretability.
+Notably, Few-shot does not necessarily improve over Zero-shot, indicating that inducing implicit generalization from a handful of examples is challenging, often failing to capture stable patterns from insufficient samples.
 5.2 How to Best Utilize Rules?
-To determine the optimal strategy for leveraging the learned probabilistic rules, we evaluated the four distinct inference methods detailed in Section 3.4.
-One might hypothesize that an LLM, with its vast pre-trained knowledge, could synthesize the rule set, their learned weights, and even a reference prediction to produce superior judgments.
-However, the empirical results, presented in Table 2, are counterintuitive.
-The most striking finding is that the simplest strategy, Linear-only, where the logistic regression model makes the final prediction, achieves the best performance on nearly all datasets.
-It outperforms all three LLM-based inference strategies in most cases, particularly in terms of F1 score where it is the top performer across all six tasks.
-This highlights the high quality of the rule set curated by RLIE and the effectiveness of the probabilistic combiner.
-Injecting more information into the LLM does not yield consistent benefits.
-Providing the LLM with rule weights (LLM + Rules + Weights) over just the rule texts does not lead to stable performance gains, suggesting that the LLM struggles to effectively internalize and apply probabilistic information to refine its reasoning.
-Most surprisingly, providing the LLM with the linear model’s own (often correct) prediction as a reference (LLM + Rules + Weights + Linear Prediction) frequently results in performance degradation compared to simply using the linear model’s prediction directly.
-This suggests that the LLM, when tasked with integrating qualitative probabilities and multiple form evidences into reasoning, can be led astray, potentially overwriting a correct judgment with an incorrect one.
-This finding aligns with existing research indicating that LLMs can be inconsistent when required to strictly adhere to complex, explicit instructions, showing the deficiency of LLMs to perform fine controlled inference.
+Table 2: Impact of inference strategies (Accuracy / Macro-F1, $\times 10^{-2}$).
+E1 is the most consistently strong approach, while prompting-based variants (E2–E4) yield mixed results.
+The best Accuracy and Macro-F1 for each backbone and dataset are bolded, respectively.
+Backbone Strategy Reviews Dreaddit Headlines Citations LLM Detect Retweets
+DeepSeek V3.2 (E1) Combiner 70.9 / 70.7 82.3 / 82.3 67.0 / 67.0 64.6 / 63.0 90.7 / 90.7 65.7 / 65.6
+(E2) Prompt(R) 65.7 / 63.1 77.5 / 76.6 66.9 / 66.8 60.4 / 56.2 89.6 / 89.6 64.4 / 64.3
+(E3) Prompt(RW) 65.4 / 64.0 77.9 / 77.1 65.2 / 65.0 55.2 / 53.5 85.2 / 85.0 62.5 / 61.8
+(E4) Prompt(RWP) 69.5 / 68.6 82.4 / 82.4 66.8 / 66.8 57.3 / 55.9 89.3 / 89.3 64.7 / 64.6
+Qwen3 235B (E1) Combiner 71.5 / 71.4 79.9 / 79.9 60.6 / 60.4 61.5 / 60.6 88.3 / 88.3 66.5 / 66.5
+(E2) Prompt(R) 64.5 / 64.2 77.3 / 76.7 60.0 / 59.7 60.4 / 60.3 88.4 / 88.4 64.5 / 63.2
+(E3) Prompt(RW) 63.4 / 63.2 77.7 / 77.1 58.5 / 58.2 60.4 / 60.1 87.1 / 87.1 65.8 / 65.0
+(E4) Prompt(RWP) 66.8 / 66.6 79.7 / 79.5 59.9 / 59.8 54.2 / 54.0 87.9 / 87.9 66.0 / 65.7
+Qwen3-Next 80B (E1) Combiner 68.3 / 67.8 81.1 / 81.1 62.9 / 62.7 56.3 / 55.0 87.6 / 87.6 61.9 / 61.8
+(E2) Prompt(R) 66.9 / 66.6 74.5 / 73.1 62.6 / 62.3 56.3 / 53.0 84.1 / 84.0 57.9 / 54.7
+(E3) Prompt(RW) 64.7 / 63.9 76.2 / 75.2 64.1 / 64.0 55.2 / 53.2 75.9 / 74.5 57.1 / 53.0
+(E4) Prompt(RWP) 67.7 / 67.4 79.5 / 79.1 62.0 / 61.9 55.2 / 53.7 87.6 / 87.6 58.9 / 56.4
+To determine the optimal inference strategy for the learned probabilistic rules, we compare the four protocols (E1–E4) defined in Section 3.4 across different LLM backbones.
+Intuitively, an LLM might synthesize the rule texts, the learned weights, and even the combiner’s reference prediction to produce superior judgments; however, Table 2 shows that this does not yield consistent gains.
+Overall, E1 (Combiner) is the most stable and consistently strong strategy: it achieves the best or near-best Accuracy and Macro-F1 in most settings, generally outperforming the three prompting-based LLM inference variants.
+This suggests that the rule set learned by RLIE is of high quality, and that an explicit probabilistic combiner can more reliably translate multi-rule evidence into final decisions.
+In contrast, injecting more information into the prompt does not reliably improve performance.
+Among all prompting-based variants, the rule-only method (E2) is often the strongest; adding rule weights (E3) does not lead to consistent improvements, suggesting that LLMs struggle to reliably exploit numeric or probabilistic signals under prompting.
+Further providing the combiner prediction as a reference (E4) is also not a dependable upgrade: while it occasionally matches or slightly exceeds E1, its performance is highly variable, rendering it unsuitable as a default strategy.
+Overall, when the model is tasked with jointly integrating multiple rules, weights, and reference signals, its final decision becomes sensitive to prompt structure and potential information conflicts, leading to inconsistent behavior.
+Taken together, our results support a robust “division of labor”: employing LLMs primarily for proposing and refining rules, while relying on the combiner for global, weight-aware aggregation and final inference.
+This synergy preserves interpretability while delivering better performance.
+We also assess the reliability of the confidence scores in Appendix C, and RLIE achieves significantly lower Expected Calibration Error compared to direct prompting in most cases, which confirms that RLIE effectively mitigates the overconfidence issue often observed in LLMs.
+5.3 Computational Cost Analysis
+We estimate the computational cost of the entire rule learning pipeline based on token usage statistics derived from our experiments (see Table 3).
+Compared to the robust baseline HypoGeniC, RLIE reduces token consumption by approximately 8%, demonstrating that our hard-example driven refinement is more sample-efficient than maintaining a large rule bank.
+Although RLIE consumes more computational resources than the lightweight IO Refinement, this additional cost is justified by our rigorous validation protocol (evaluating rules on the full validation set vs.
+only 10 samples in IO Refinement), which is crucial for ensuring the reliability of the learned neuro-symbolic system.
+Table 3: Average computational cost per task.
+RLIE achieves a balance between efficiency and robustness, consuming fewer tokens than HypoGeniC while performing comprehensive validation.
+Energy costs are estimated by assuming a conservative energy intensity of 1.0 J/token for the MoE backbone.
+Method Avg. Tokens (M) Est. Energy (kWh) Relative Cost
+HypoGeniC 13.24 3.68 108%
+RLIE (Ours) 12.23 3.40 100%
+IO Refinement 08.45 2.35 069%
+5.4 Case Study: Rule Evolution and Representation
+Table 4: Iterative Rule Refinement Process (Retweet Task).
+RLIE refines the rule bank in an error-driven manner.
+At round $r=0$, we start from a randomly sampled example; at round $r>0$, we sample a hard example misclassified by the round r1r\!-\!1 model.
+[NEW]  newly added rules, [KEEP]  inherited active rules, and [DROP]  rules with zeroed weights.
+Text in bold indicates the semantic focus.
+Task: Predict which of two tweets is retweeted more. Output: First vs Second (0/1).
+Rd Prompt (Side-by-side) Rule Bank Update (Full Text) Metric
+0 T1: … #House debates issue of #health insurance policies eliminated
+T2: … #House voting on bill to restore #health insurance policies …
+GT: Second (1)
+[NEW] 1. Tweets that use stronger emotional language or dramatic framing …are more likely to be retweeted. (w=.035)
+[NEW]
+ 2. Tweets that include vivid details, quotes, or specific examples outperform those that offer general summaries. (w=.008)
+[DROP]
+[…]
+Acc: 0.625
+F1: 0.625
+1 T1: Andrew C. McCarthy: The Scheme behind the Obamacare Fraud …
+T2: Like all swindles, Obamacare cannot work if its targeted victims …
+GT: Second (1); Prev: First (0)
+[NEW] 1. Tweets that prioritize clarity and immediate comprehension …are favored in news and public-affairs contexts. (w=.045)
+[KEEP]
+ 2. Tweets that use stronger emotional language or dramatic framing …are more likely to be retweeted. (w=.034)
+[DROP]
+[…]
+Acc: 0.680
+F1: 0.679
+2 T1: NEW VID!! :D ***MEET MY DAD!*** … RT ? :)
+T2: what do you think about the SNEAK PEEK … ***MEET MY DAD!*** …
+GT: Second (1); Prev: First (0)
+[NEW] 1. Tweets that use personal voice or self-referential framing …are more retweetable when they convey authenticity. (w=.223)
+[KEEP]
+ 2. Tweets that prioritize clarity and immediate comprehension …are favored in news contexts. (w=.178)
+[DROP]
+[…]
+Acc: 0.700
+F1: 0.699
+To investigate the inner workings of RLIE, we conduct two case studies: one tracing the dynamic trajectory of rule refinement on Retweets, and another analyzing the semantic efficiency of the final rule set on Headlines.
+Dynamics of Rule Evolution.
+We track the iterative updates of the rule set in Table 4.
+The process acts as a continuous error-correction loop.
+Initial heuristics like urgency or novelty start with non-zero weights but are effectively pruned to zero as they fail to generalize.
+Meanwhile, hard examples trigger the generation of more specific semantic features, such as clarity in news contexts, which are then up-weighted.
+This confirms that RLIE actively performs soft feature selection, evolving from coarse heuristics to robust, task-aligned discriminators.
+Table 5: The rule set learned by RLIE.
+The logistic combiner assigns high importance to Surprise and Relatability features while effectively pruning the Urgency rule ($w=0$) via $L_{1}$ regularization.
+Rule ID Wgt. Learned Rules Focus
+Rule 2 0.039 Headlines that frame content as a surprising outcome… provoke curiosity. Surprise
+Rule 0 0.005 Headlines that frame content as a direct, relatable scenario Relatable
+Rule 1 0.000 Headlines that incorporate a sense of urgency (Pruned)
+Figure 3: t-SNE of rule embeddings. RLIE learns complementary rules while HypoGeniC shows semantic noise and IO Refinement exhibits mode collapse (visualizing all refinement steps).
+Efficiency of Representation.
+We visualize the embeddings of the final rule sets for the Headlines task using t-SNE in Figure 3 and inspect the learned rules in Table 5.
+HypoGeniC exhibits a scattershot distribution, retaining 20 diverse but noisy rules that yield suboptimal accuracy.
+IO Refinement suffers from mode collapse, where all 5 retained rules cluster tightly and redundantly focus on aggressive imagery.
+In contrast, RLIE achieves semantic efficiency, converging to just three widely separated rules that capture orthogonal concepts: Surprise and Relatability.
+Notably, the Urgency rule is correctly identified as ineffective and pruned, mirroring the trajectory seen in the Retweets case.
+This disentanglement allows RLIE to achieve the highest accuracy with the most compact representation.
 6 Discussion
-Our hierarchical evaluation reveals a central finding: achieving stable, globally consistent reasoning by prompting an LLM with a set of explicitly weighted rules is non-trivial.
-The simplest “Linear-only” strategy, which uses the learned logistic regression model as the final classifier, performed best on nearly all datasets and achieved the highest F1 score across the board.
-In contrast, injecting the rule texts, their weights, and even the linear model’s reference prediction back into the LLM often failed to yield stable improvements and, in some cases, degraded performance.
-This result aligns with the observation that LLMs excel at semantic generation and interpretation but are less reliable at fine-grained, controlled probabilistic integration.
-This suggests that a more robust paradigm involves a clear division of labor: leveraging LLMs for local, semantic tasks such as judging individual rules and generating new candidates, while entrusting a transparent and calibratable probabilistic combiner—in our work, a logistic regression model with elastic net regularization—with the global tasks of weighting and selection.
-From a broader perspective, using natural language rules as the fundamental unit of reasoning bypasses the rigid dependency on predefined predicate spaces found in traditional methods, offering greater expressiveness for open-ended semantics and unstructured data.
-At the same time, this approach avoids overwhelming the LLM’s context with complex probabilistic details from an entire rule set, thereby preserving its ability to make stable judgments at the single-rule level.
-Unlike frameworks that model interactions at the predicate level (e.g., Markov Logic Networks), we advocate for enhancing the capabilities of the global combiner while keeping the local interface—the LLM’s ternary judgments and coverage information—fixed.
-This linear log-odds combiner can be naturally extended to more sophisticated models, such as (i) interpretable additive models (e.g., GAMs) to capture rule interactions; (ii) message passing on a factor graph to globally harmonize rule confidences; (iii) Bayesian logistic regression with hierarchical priors (e.g., ARD, Laplace) for joint uncertainty and sparsity modeling; or (iv) techniques like Platt scaling or isotonic regression to explicitly calibrate the LLM’s ”abstain” outputs.
-These extensions are fully compatible with the current RLIE architecture.
-Overall, we propose a replicable engineering principle: LLMs are responsible for semantic generation and local judgment, while traditional probabilistic learning methods handle global aggregation and uncertainty management.
-On this foundation, stronger graphical or additive structures and more advanced inference algorithms can be progressively introduced to build more flexible and powerful, yet still interpretable and calibratable, neuro-symbolic systems.
+We propose a general paradigm for designing inference systems where natural-language rules serve as the basic unit of reasoning.
+In this setting, future improvements may stem less from injecting global probabilistic details into LLM prompting, and more from shaping a reusable interface that separates LLM-based semantics from probabilistic learning.
+Compared to traditional predicate-based frameworks, natural-language rules better accommodate open-ended semantics and unstructured data; however, asking an LLM to simultaneously interpret rules and perform global weighted aggregation can be sensitive to prompt phrasing and instruction competition, potentially hurting stability.
+A promising direction is to fix the local interface—letting the LLM provide per-rule signals (e.g., judgments or confidence scores) alongside coverage information—and to delegate global consistency, conflict resolution, and uncertainty management to a non-LLM, transparent aggregation module.
+This module can be further strengthened by interpretable additive models or explicit calibration techniques, enabling robust neuro-symbolic systems that combine the semantic flexibility of LLMs with the rigor of classical statistical models.
 7 Conclusion
-In this paper, we addressed the critical challenge of learning robust and interpretable rule sets from data.
-While Large Language Models (LLMs) have created new possibilities for generating rules in natural language, existing methods have struggled with the principled aggregation and calibration of these rules.
-We introduced RLIE, a unified framework that synergizes the rule-generation capabilities of LLMs with the global optimization of probabilistic models.
-By integrating rule generation, regularized logistic regression, and error-driven iterative refinement, RLIE produces a compact and calibrated set of weighted, natural language rules.
-Our experiments demonstrate that this hybrid approach is highly effective.
-However, directly using LLMs to apply weighted rules degrades the performance of downstream tasks, which further unveils the limitations of applying LLMs in rule based reasoning.
-Our work underscores the value of a neuro-symbolic division of labor, where LLMs handle local semantic tasks and classical models manage global probabilistic reasoning.
-The paper investigates the potentials and limitations of using LLMs for inductive reasoning tasks, proposing a unified framework which integrates LLMs with classic probabilistic rule combination methods, paving the way for more building reliable AI and contribute to knowledge discovery.
-8 Ethic Statement
-This research adheres to the ICLR Code of Ethics.
-The datasets used in our experiments are derived from the HypoBench benchmark, which is composed of publicly available data from prior academic research.
-To the best of our knowledge, personally identifiable information has been anonymized in these datasets.
-We acknowledge that the rules generated by our method reflect the patterns and potential societal biases present in the underlying training data.
-While a primary goal of our work is to enhance interpretability, this also means that any learned biases will be transparent.
-We caution that the application of such rule-based systems, particularly in sensitive domains like mental health analysis or persuasion, should be accompanied by careful human oversight to mitigate the risk of misuse or the perpetuation of harmful stereotypes.
-Our work is intended to contribute to the development of more transparent and auditable AI systems.
-9 Reproducibility Statement
-We are committed to ensuring the reproducibility of our research.
-The complete source code for the RLIE framework, along with the scripts for running all experiments, will be made publicly available upon publication.
-All datasets used in this study are from the public HypoBench benchmark, and we will provide the exact data splits to facilitate direct comparison.
-A detailed description of our experimental setup, including model specifications, baseline implementations, and primary hyperparameters, is provided in Section 4.
-A comprehensive list of all prompts used to interact with the LLM for rule generation, judgment, and inference is included in the Appendix C.
-The code will be released once the paper is accepted.
+In this work, we introduced RLIE, a unified framework for learning compact, interpretable natural-language rule sets by integrating LLM-based hypothesis generation with a calibrated probabilistic combiner and error-driven refinement.
+Extensive experiments across diverse real-world tasks and LLM backbones demonstrate that direct inference via the learned logistic combiner yields the most robust and consistent performance, whereas re-injecting rules and weights into LLM prompting often leads to mixed results or degradation.
+These findings advocate for a principled neuro-symbolic division of labor: LLMs excel at local semantic operations—such as proposing hypotheses and judging applicability—while classical probabilistic models are superior for global selection, weighting, and calibration.
+This synergy enables the construction of reliable, auditable, and high-performance reasoning systems in natural language.
+Impact Statement
+This work advances machine learning by studying how LLMs can generate natural-language decision rules and how classic probabilistic modeling can combine these rules into calibrated, inspectable predictions.
+A key positive impact is improved transparency and controllability: explicit rule sets and learned weights make it easier to audit model behavior, attribute errors to specific rules, and iteratively refine systems to better match desired policies, which can be valuable for text classification and screening workflows where stakeholders require understandable rationales.
+The framework may also lower the barrier to building lightweight, modular decision systems by reusing and updating rule sets without retraining large models end-to-end.
+As with other data-driven approaches, risks include learning rules that encode biases or spurious correlations present in data, or over-trusting rule-based rationales in high-stakes settings.
+These risks can be mitigated through careful dataset curation and documentation, evaluation across subpopulations and robustness tests, and maintaining human oversight when deploying or updating rule sets.
 Appendix A More Details
@@
 The primary challenge lies in identifying subtle linguistic cues, such as exaggerated sentiment or vague descriptions, which are difficult for even human evaluators to discern reliably.
-Persuasive Argument Prediction (Dreddit).
-Given a pair of argumentative texts, the model must predict which is more persuasive.
-This task emphasizes linguistic features correlated with persuasiveness, such as argument structure, rhetorical devices, and emotional tone.
-It is particularly challenging as persuasiveness often depends on complex context and psychological factors.
+Persuasive Argument Prediction (Dreaddit).
+The objective is to detect mental stress signals from Reddit posts across different communities.
+These posts, sourced from social media, often contain complex personal narratives that reflect an individual’s psychological state.
+The primary challenge lies in investigating specific linguistic features indicative of mental stress, identifying subtle patterns in the text that effectively signal the presence of distress.
 News Headline Engagement (Headlines).
@@
 Due to the highly stochastic nature of social media propagation, this task places stringent demands on the explanatory power and stability of the learned rules.
-Appendix B LLM Usage
+A.4 Hyperparameter Settings
+Iterative Refinement.
+To balance performance and computational cost, we set the maximum number of refinement iterations $R_{max}=20$.
+The early stopping mechanism uses a patience of $p=3$ rounds with a minimum improvement threshold $\delta=1\times 10^{-3}$ on the validation set accuracy.
+Logistic Regression.
+We implement the logistic combiner using the scikit-learn library.
+Hyperparameters for Elastic Net are selected via 5-fold stratified cross-validation on $\mathcal{S}_{train}$.
+We perform a grid search for the inverse regularization strength $C$ (30 logarithmically spaced values in $10^{-3},10^{1}]$) and the L1 ratio $\alpha$ (30 linearly spaced values in $[0.01,0.99]$).
+LLM Generation.
+For the Rule Generation and Refinement stages, we prioritize stability and reproducibility.
+We set the generation temperature to $1\times 10^{-5}$.
+The sampling process uses a batch size of 20 training samples, generating 5 rules per API call, with a maximum token limit of 8000.
+Appendix B Parameter Study
+Table 6: Sensitivity analysis of the coverage threshold $\gamma$ on the Headline dataset (Backbone: DeepSeek-V3). Performance remains robust for $\gamma\in[0.1,0.5]$.
+Coverage Threshold ($\gamma$) Accuracy F1 Score
+0.1 66.7 66.7
+0.2 67.0 67.0
+0.3 67.3 67.2
+0.4 66.5 66.5
+0.5 66.5 66.5
+0.6 65.5 65.5
+0.7 66.5 66.5
+0.8 65.6 65.6
+0.9 64.3 64.2
+Appendix C Calibration Analysis
+To validate our claim that the logistic combiner provides calibrated probabilistic outputs, we evaluate the Expected Calibration Error (ECE) (Guo et al., 2017) of RLIE compared to the rule-injection prompting baseline (E2).
+C.1 Methodology
+We compute ECE using 10 equal-width bins.
+For both methods, ECE measures the weighted average difference between the predicted confidence and the empirical accuracy: $\mathrm{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{N}|\mathrm{acc}(B_{m})-\mathrm{conf}(B_{m})|$.
+  • RLIE (Ours): The probability is directly derived from the logistic regression combiner: $p(y=1|x)=\sigma(\beta_{0}+\sum w_{j}r_{j}(x))$.
+Since logistic regression minimizes the negative log-likelihood, it naturally aligns predicted probabilities with empirical frequencies.
+  • Prompting (E2): Since standard prompting yields binary text, we adapt it to produce probabilistic outputs for fair comparison.
+We constrain the LLM (DeepSeek-V3.2) to output a single token (e.g., ’A’ or ’B’) and extract the token log-probabilities.
+The probability of the positive class is computed via softmax: $p=\frac{\exp(\ell_{pos})}{\exp(\ell_{pos})+\exp(\ell_{neg})}$.
+C.2 Results
+Table 7 reports the mean ECE and standard deviation over 3 runs.
+RLIE achieves consistently lower ECE scores (better calibration) on 5 out of 6 datasets.
+Notably, on tasks like Retweets and Reviews, Prompting exhibits high calibration error (>0.45>0.45), indicating severe overconfidence (predictions are often close to 1.0 even when wrong).
+In contrast, RLIE maintains an ECE around 0.11, demonstrating that the learned weights serve as reliable uncertainty estimates.
+Table 7: Comparison of Expected Calibration Error (ECE) on the test set (Lower is better). RLIE significantly outperforms Prompting (E2) on most tasks, providing more reliable confidence estimates.
+Dataset RLIE (Ours) Prompting (E2)
+ECE $\downarrow$ ECE $\downarrow$
+Retweets 0.116 $\pm$ 0.041 0.480 $\pm$ 0.040
+Reviews 0.113 $\pm$ 0.022 0.511 $\pm$ 0.065
+Headlines 0.104 $\pm$ 0.029 0.371 $\pm$ 0.034
+Dreaddit 0.226 $\pm$ 0.054 0.368 $\pm$ 0.050
+Citations 0.179 $\pm$ 0.048 0.494 $\pm$ 0.261
+LLM Detect 0.252 $\pm$ 0.052 0.187 $\pm$ 0.051
+Appendix D LLM Usage
 Large Language Models (LLMs) were used as an assistive tool in preparing this manuscript.
@@
 The authors have reviewed and edited all LLM-generated content and take full responsibility for the final manuscript.
-Appendix C Prompts
+Appendix E Prompts
 Here we are providing the prompts we are using to conduct all the experiments.